### v1

# Functional Liftings of Restricted Geometric Inequalities

Abstract.
We investigate what we term “generalized sup-convolutions”. We show that functional inequalities that enjoy an interpretation as sup-convolution inequalities can be deduced from the special case of indicator functions corresponding to a geometric inequality. As consequences we derive a Borell-Brascamp Lieb inequality for the Gaussian Brunn-Minkowski inequality and give a functional analog of the log-Brunn-Minkowski conjecture. We also show that Barthe’s reverse Brascamp-Lieb inequality is equivalent to the geometric inequality it implies, and give a functional formulation of a conjecture of Schneider. Though we focus on Euclidean applications, our results are general and can be directly applied in more abstract settings, like groups or even topological measure spaces without algebraic structure, we instantiate this claim with a Borell-Brascamp-Lieb type inequality for nilpotent Lie groups.
1. Introduction
In this article we investigate the equivalence of certain types of geometric inequalities and their functional analogs, each of which can be reviewed as reverse Hölder type inequalities. Our setting will be abstract and rather flexible, but let us start with an example: the well-known equivalence of the celebrated Brunn-Minkowski inequality (see [66]), a linchpin for modern convex geometry, and the Prékopa-Leindler inequality (see [59, 38]), it’s functional version possessing broad applications in various fields throughout mathematics.
The Brunn-Minkowski inequality asserts that, for any pair of Borel sets $A,B\subset{\mathbb{R}}^{n}$, the following volume inequality
(1) $|A+B|^{\frac{1}{n}}\geq|A|^{\frac{1}{n}}+|B|^{\frac{1}{n}},$
where $|\cdot|$ (or $|\cdot|_{n}$) is the $n$-dimensional Lebesgue volume, is valid. The Prékopa-Leindler inequality asserts that for any triple of non-negative measurable functions $f,g$ and $h$ such that
(2) $h((1-t)x+ty)\geq f^{1-t}(x)g^{t}(y),\qquad\mbox{for some }t\in[0,1]\mbox{ and all }x,y\in\mathbb{R}^{n}$
the integrated inequality
(3) $\int h\geq\left(\int f\right)^{1-t}\left(\int g\right)^{t}$
holds.
The numerous extensions and the ramifications of the Brunn-Minkowski and Prekopa-Leindler inequality has had in various fields of mathematics are detailed in the survey of Gardner [27]. An exhaustive list of references is beyond the scope of this paper. We would like to highlight two very important and classical papers of Borell [10], and Brascamp and Lieb [15], where generalizations of both of these inequalities have appeared; such inequalities have become referred to as “Borell-Brascamp-Lieb” and “Brascamp-Lieb” inequalities. More recent examples of such works, as well as modern treatments, can be found in the books [3, 4, 66].
We now return to our original point of investigation: the equivalence of the inequalities (1) and (3).
Taking $f=\mathbbm{1}_{A}$ and $g=\mathbbm{1}_{B}$, in (3), yields
(4) $|(1-t)A+tB|\geq|A|^{1-t}|B|^{t}$
which, after a clever choice of $t$, recovers (1). The converse direction can be proven by showing that Brunn-Minkowski implies Prékopa-Leindler in one-dimension and then using the tensorization property of (3) to recover the $n$-dimensional functional inequality, or alternatively interpreting the integrals in (3) as the volumes of $(n+1)$-dimensional sets.
However, in more general settings when a geometric inequality holds only for a restricted class of sets and/or a particular measure on a vector space or group, it can be less clear how to implement such approaches. For example, the conjecture of Gardner and Zvavitch [30] asked if
(5) $\gamma((1-t)A+tB)\geq\left[(1-t)\gamma(A)^{\frac{1}{n}}+t\gamma(B)^{\frac{1}{n}}\right]^{n}$
held for $A,B$ origin-symmetric convex bodies in ${\mathbb{R}}^{n}$ (see Section 6). By Borell [10], an $n$-dimensional probability measure that satisfies (5) for all $A,B$ Borel sets is necessarily the uniform measure on a bounded convex set, thus such an inequality necessarily fails for $\gamma$ and general $A,B$. Nevertheless this question was answered in the affirmative by Eskenazis and Moschidis [24]. Our first result is a functional version of (5). It says that with the additional assumption that the relevant functions are even and unimodal [^1] We consider a function $f\colon{\mathbb{R}}^{n}\to{\mathbb{R}}$ unimodal if its super-level sets are convex bodies (see Section 2), the Borell-Brascamp-Lieb inequality of the Lebesgue measure holds for the Gaussian.
Theorem 1.1.
Let $\gamma$ be the standard Gaussian probability measure on ${\mathbb{R}}^{n}$, let $t\in[0,1]$, $\alpha\geq-\frac{1}{n}$, let $f,g\colon{\mathbb{R}}^{n}\to[0,\infty)$ be even, unimodal functions. Assume that $h\colon{\mathbb{R}}^{n}\to[0,\infty)$ is any Borel measurable function for which
$h((1-t)x+ty)\geq[(1-t)f(x)^{\alpha}+tg(y)^{\alpha}]^{\frac{1}{\alpha}}$
holds for all $x,y\in{\mathbb{R}}^{n}$. Then
$\int_{{\mathbb{R}}^{n}}hd\gamma\geq\left[(1-t)\left(\int_{{\mathbb{R}}^{n}}fd\gamma\right)^{\beta}+t\left(\int_{{\mathbb{R}}^{n}}gd\gamma\right)^{\beta}\right]^{\frac{1}{\beta}},$
where $\beta=\frac{\alpha}{1+n\alpha}$. Moreover, this functional inequality is equivalent to (5).
We mention that [1, Theorem 1.7] is a special case of Theorem 1.1 when $f,g$ are assumed to be $\log$-concave and $\alpha\geq 0,$ which in turn is the Gaussian special case of Theorem 3.3 to come. The $\alpha=0$ case corresponds to (2) and (3). The key to our approach is the study of properties held by what we will refer to as a sup-convolution [^2] Note that this a multiplicative version of a sup-convolution, when $f=e^{-V},g=e^{-W}$, $\sup_{(1-t)x+ty=z}f^{1-t}(x)g^{t}(y)=\exp\left\{-\inf_{(1-t)x+ty=z}\bigg{[}(1-t)V(x)+tW(y)\bigg{]}\right\}.$ The quantity $\inf_{(1-t)x+ty=z}\bigg{[}(1-t)V(x)+tW(y)\bigg{]}$ is referred to as the inf-convolution of $V$ and $W$. . For instance in the $\alpha=0$ case, given a pair of non-negative measurable functions $f,g\colon{\mathbb{R}}^{n}\to[0,\infty)$, we set
[…]
That is, $f\square g$ is the smallest function that satisfies (2). Indeed, any $h$ satisfying (2), satisfies $h(z)\geq f\square g(z)\geq f^{1-t}(x)g^{t}(y)$ for (1t)x+ty=z(1-t)x+ty=z, so that (neglecting potential measurability issues) we have
$\int h\geq\int f\square g\geq\left(\int f\right)^{1-t}\left(\int g\right)^{t}.$
A property of the $\square$ operation that proves pivotal, beyond some reasonably expected monotonocity and measurability (see Section 2 for a complete development), is a somewhat peculiar, restricted reverse triangle inequality. That is given two pairs of functions $(f_{0},g_{0})$ and $(f_{1},g_{1})$ such that $f_{1}(x)>0$ implies that $f_{0}(x)\geq f_{0}(y)$ for all $y\in\mathbb{R}^{n}$, and $g_{1}(x)>0$ implies $g_{0}(x)\geq g_{0}(y)$ for all $y\in\mathbb{R}^{n}$, then
$(f_{0}+f_{1})\square(g_{0}+g_{1})\geq f_{0}\square g_{0}+f_{1}\square g_{1}.$
We will see that the same behavior arises in various circumstances, for instance an analogous approach will yield the following result for the geometric mean of convex bodies (see Section 5 for further details).
Theorem 1.2.
Let $t\in[0,1]$. Then the Lebesgue measure satisfies
$|A^{1-t}B^{t}|\geq|A|^{1-t}|B|^{t}$
for all even convex bodies $A$ and $B$, if and only if $f,g\colon{\mathbb{R}}^{n}\to[0,\infty)$ even and unimodal functions, such that
$h(z)\geq f(u)^{1-t}g(v)^{t},$
for $u,v\in C(z;f,g)$, where
$C(z;f,g)\coloneqq\{(u,v)\colon z\in\{f\geq f(u)\}^{1-t}\{g\geq g(v)\}^{t}\}.$
implies
$\int_{\mathbb{R}^{n}}h\geq\left(\int_{\mathbb{R}^{n}}f\right)^{1-t}\left(\int_{\mathbb{R}^{n}}g\right)^{t}.$
That is, Theorem 1.2 gives an equivalent functional version of the celebrated $\log$-Brunn-Minkowski conjecture of Böröczky, Lutwak, Yang, and Zhang [14]. In particular, Theorem 1.2 affirmatively answers a question posed by Andrea Colesanti “Is there a functional version of the log-Brunn-Minkowski inequality?” at the conference 61 Probability Encounters, in honour of Sergey Bobkov held at Paul Sabatier University, Toulouse, France in the summer of 2023 [17, Main Question].
Moreover, in the unconditional case [^3] A function $f\colon{\mathbb{R}}^{n}\to{\mathbb{R}}$ is said to be unconditional if, for each $\varepsilon_{i}\in\{-1,1\}$ $i=1,\dots,n$, and $x=(x_{1},\dots,x_{n})\in{\mathbb{R}}^{n}$, $f(\varepsilon_{1}x_{1},\dots,\varepsilon_{n}x_{n})=f(x)$; a set is unconditional if its characteristic function is an unconditional function., where the log-Brunn-Minkowski conjecture has been verified (see [63]), we have the following result.
Theorem 1.3.
Let $t\in[0,1]$. For $f=(f_{1},f_{2})$, with $f_{1},f_{2}\colon{\mathbb{R}}^{n}\to[0,\infty)$ even, unconditional, and unimodal functions, and suppose that $h$ satisfies
$h(z)\geq f_{1}(u)^{1-t}f_{2}(v)^{t},$
for $(u,v)$ such that $z\in\{f_{1}\geq f_{1}(u)\}^{1-t}\{f_{2}\geq f_{2}(v)\}^{t}$. Then the $\log$-Prékopa-Leindler inequality holds:
(7) $\int_{{\mathbb{R}}^{n}}h\geq\left(\int_{{\mathbb{R}}^{n}}f_{1}\right)^{1-t}\left(\int_{{\mathbb{R}}^{n}}f_{2}\right)^{t}.$
In fact, all of the above results can be considered corollaries of Theorem 2.3 below. Among other implications to come is the following result on the saturation of Barthe’s reverse Brascamp-Lieb inequality by the geometric inequality it implies.
Theorem 1.4.
Let $n,m$ be positive integers. Consider integers $1\leq n_{1},\dots,n_{m}<n$ and constants $c_{1},\dots,c_{m}>0$ such that $\sum_{i}c_{i}n_{i}=n$, and let $c=\sum_{i}c_{i}$, and for each $i$, set $t_{i}=\frac{c_{i}}{c}$. For each $i=1,\dots,m$, assume that $B_{i}:{\mathbb{R}}^{n}\to{\mathbb{R}}^{n_{i}}$ is a surjective linear map, and that $\cap_{i}\text{ker}(B_{i})=\{0\}$.
Barthe’s reverse Brascamp-Lieb inequality (see Inequality (10) below) is equivalent to the following geometric inequality: for any collection of Borel sets $A_{1},\dots,A_{m}$, with $A_{i}\subset{\mathbb{R}}^{n_{i}}$, we have
$\left|\sum_{i=1}^{m}c_{i}B_{i}^{*}A_{i}\right|\geq E\mathcal{M}_{0}^{(c)}(|A_{i}|_{n_{i}}).$
We refer the reader to Section 2, Equation (8) for the definition of $\mathcal{M}_{0}^{(c)}$ and to Section 4, Equation (11), for that of the constant $E$. and note that this saturation is anticipated by the recent work of [13] who have already shown the set theoretic case to be extremal in the special case of “geometric instances”.
Further we show the equivalence of another conjecture and a functional analog. More explicitly Schneider conjectured in [65] that among convex bodies of a fixed volume, the volume of the $m$-th order difference body $D^{m}(K)$ (see Conjecture 4.4) is minimized by ellipsoids when $n\geq 3$ and $m\geq 2$. We show in Section 4 that this is equivalent to the following functional inequality for $f:\mathbb{R}^{n}\to[0,\infty)$ unimodal,
$\int_{(\mathbb{R}^{m})^{n}}\left(\sup_{x\in\mathbb{R}^{n}}\prod_{i=0}^{m}f(x-z_{i})\right)^{\frac{1}{m+1}}dz\geq c(m,n)\left(\int_{\mathbb{R}^{n}}f^{\frac{1}{m}}\right)^{m},$
where $c(m,n)\coloneqq|D^{m}(B)|_{nm}|B|^{m}$ and $B$ is the $\ell_{2}$ unit ball. In the same article, Schneider proved that the volume of the $m$-th order difference body for $n=2$ and any $m$ is minimized by centrally symmetric convex bodies.
As mentioned in the abstract the techniques developed are not limited to Euclidean spaces. In fact, the ideas here are inspired by, and could be used to recover, the proof of a Prekopa-Leindler type inequality on the integers, by the second author and Marsiglietti [55, Theorem 3.8]. To further hint at broader application, we also give a Borell-Brascamp-Lieb inequality for nilpotent groups, generalizing Tao [67], which in turn generalized the case given for the Heisenberg group [56].
Theorem 1.5.
Let $G$ be a simply connected nilpotent Lie group of (topological) dimension ddwith a Haar measure $\mu$ and $\mathcal{M}_{p}^{(t)}(u,v)$ the $t$-weighted $p$-mean [^4] We refer to the next section and more precisely to (8) for the definition of $\mathcal{M}_{p}^{(t)}$. of $u$ and $v$. Then, for $\alpha\in[-\frac{1}{d},1]$ and $t\in(0,1)$, if
$h(xy)\geq\mathcal{M}_{\alpha}^{(t)}(f(x),g(y))$
holds for all $x,y$, then
$\int_{G}h\ d\mu\geq\mathcal{M}_{\alpha^{\prime}}^{(t)}\left(\frac{\int_{G}f\ d\mu}{(1-t)^{d}},\frac{\int_{G}g\ d\mu}{t^{d}}\right).$
where $\alpha^{\prime}=\frac{\alpha}{1+\alpha d}$.
Let us outline the remainder of the article. In Section 2 we introduce an abstract sup-convolution, and derive our main technical tool in Theorem 2.3. In Section 3 we derive the dimensional Brunn-Minkowski inequality for the Gaussian measure and generalizations thereof. In Section 4 we deal with the equivalences betweeen set theoretic and function theoretic formulations of Barthe’s reverse Brascamp-Lieb inequality and Schnieder’s conjecture. In Section 5 we give a functional version of the log-Brunn-Minkowski conjecture and related results for the $L_{p}$-Brunn-Minkowski theory. Finally in Section 6 we discuss further generalizations of the development in Section 2, limitations of our methods, and give a Borell-Brascamp-Lieb type inequality for nilpotent Lie groups.
2. Generalized Sup-Convolutions
In this section we introduce some well-known objects from convex geometry (convex body, means) and the new notion of generalized or abstract sup-convolution (Definition 2.2) and prove two related abstract theorems that constitute the starting point of our investigations.
A set $K\subseteq\mathbb{R}^{n}$ is a body, if $K$ is compact with non-empty interior. $K$ is a convex body, if it is body that satisfies $(1-t)K+tK\subseteq K$ for $t\in[0,1]$. We call a function $f:\mathbb{R}^{n}\to\mathbb{R}$ unimodal if
$\{f\geq t\}=\{x\in\mathbb{R}^{n}:f(x)\geq t\}$
is a convex body for $t\in\mathbb{R}$ when it is non-trivial (that is non-empty and not all of $\mathbb{R}^{n}$). Many of the results to come will follow with lighter assumptions, for instance, compactness can often be replaced by Borel measurability, but we will not pursue these technicalities.
For $u=(u_{1},\dots,u_{n})$ with $u_{i}>0$, $t=(t_{1},\dots,t_{n})$ with $t_{i}>0$ and $p\in(-\infty,0)\cup(0,\infty)$ we define the “mean” ($\sum_{i}t_{i}$ is not necesarily set to be one)
(8) $\mathcal{M}_{p}^{(t)}(u)=\left(\sum_{i=1}^{n}t_{i}u_{i}^{p}\right)^{\frac{1}{p}}.$
For $p\in\{-\infty,0,\infty\}$ we define
$\mathcal{M}_{-\infty}^{(t)}(u)\coloneqq\min_{i}\{u_{i}\},\qquad\qquad\mathcal{M}_{0}^{(t)}(u)\coloneqq\prod_{i=1}^{n}u_{i}^{t_{i}},\qquad\qquad\mathcal{M}_{\infty}^{(t)}(u)\coloneq\max_{i}\{u_{i}\}.$
Note that for $p\in\{\pm\infty\}$ this is the continuous extension of $\mathcal{M}_{p}^{(t)}$; while under the assumption that $\sum_{i=1}^{n}t_{i}=1$ it is the continuous extension for $p=0$ as well. We further extend $\mathcal{M}_{p}^{(t)}$ to $0,\infty)^{n}$ by
$\mathcal{M}_{p}^{(t)}(u)=0$
for $u$ such that $\prod_{i=1}^{n}u_{i}=0$.
For later purpose, let us observe that Hölder’s inequality gives
(9) $(p+q)\mathcal{M}_{p}^{(t)}(u)\mathcal{M}_{q}^{(t)}(v)\geq(p+q)\mathcal{M}_{r}^{(t)}(uv),$
where $(uv)_{i}=u_{i}v_{i}$, for $p,q,r\in\mathbb{R}$, satisfying $\frac{1}{r}=\frac{1}{p}+\frac{1}{q}$. Note, all of the inequalities are trivial if either $u$ or $v$ have 0 in any coordinate, and that $\frac{1}{r}=\frac{1}{p}+\frac{1}{q}$ has no solution in $r$ when $p+q=0$. However, when $p+q>0$ we have
$\mathcal{M}_{p}^{(t)}(u)\mathcal{M}_{q}^{(t)}(v)\geq\mathcal{M}_{r}^{(t)}(uv)$
and taking $q>0$ and $p\downarrow-q$, so that $r\downarrow$-$\infty$ we have by continuity,
$\mathcal{M}_{-q}^{(t)}(u)\mathcal{M}_{q}^{(t)}(v)\geq\mathcal{M}_{-\infty}^{(t)}(uv).$
This last fact is also a consequence of the inequality $\sum_{i}t_{i}x_{i}\leq C\sum_{i}t_{i}y_{i}$, for positive numbers $x_{i}\leq Cy_{i}$.
Given a collection of sets, the next definition constructs a corresponding collection of functions.
Definition 2.1.
Given a Polish measure space $(E,\mathscr{E},\mu)$ and a collection of subsets $\mathcal{A}\subseteq\mathscr{E}$ containing both the empty set, we let $\mathcal{F}(\mathcal{A})$ denote the collection of functions
$\mathcal{F}(\mathcal{A})\coloneqq\{f:E\to\mathbb{R}\ |\ \{f\geq t\}\in\mathcal{A}\hbox{ for all }t\in\mathbb{R}\}.$
Hereafter, we will assume that a collection of sets $\mathcal{A}$ contains the empty set, even if it is not explicitly mentioned. For example, we may say, let $\mathcal{A}$ be the collection of origin symmetric convex bodies, in place of let $\mathcal{A}$ be the collection of origin symmetric convex bodies and the empty set.
For $p>0$ and $f\in\mathcal{F}(\mathcal{A})$ we denote $\|f\|_{p}\coloneqq\left(\int|f|^{p}d\mu\right)^{\frac{1}{p}}$ and emphasize that the underlying measure $\mu$ is understood.
Let $\{(E_{1},\mathscr{E}_{1},\mu_{1}),\dots,(E_{n},\mathscr{E}_{n},\mu_{n})\}$ and $(F,\mathscr{F},\nu)$ be measure spaces. We will denote $E=E_{1}\times\cdots\times E_{n}$ and $\mu=\mu_{1}\otimes\cdots\otimes\mu_{n}$. For $f=(f_{1},\dots,f_{n})$ with $f_{i}:E_{i}\to\mathbb{R}$, and $x\in E$, set
$f(x)\coloneqq(f_{1}(x_{1}),\dots,f_{n}(x_{n})).$
For vectors $u,v\in\mathbb{R}^{n}$ we write $u\leq v$ when $u_{i}\leq v_{i}$ for all $i$. We assume that there exists on each $E_{i}$ a privileged class of sets $\mathcal{A}_{i}$, and we will consider a generalized sup-convolution of $f_{i}\in\mathcal{F}(\mathcal{A}_{i})$.
We are now in position to introduce the generalized sup-convolution that constitutes one of the central objects of this paper.
Definition 2.2.
We consider a functional $\square$ that takes a collection $f=(f_{1},\dots,f_{n})$ of non-negative $f_{i}\in\mathcal{F}(\mathcal{A}_{i})$ to $\square f:F\to[0,\infty]$ to be a generalized sup-convolution if the following are satisfied for $f_{i},g_{i}\in\mathcal{F}(\mathcal{A}_{i})$:
  1. (1)
    $f_{i}\leq g_{i}$ for all $i$ implies $\square f\leq\square g$
  2. (2)
    If each $g_{i}$ has the property that $g_{i}(x)>0$ implies then $f_{i}(x)\geq f_{i}(y)$ for all $y\in E_{i}$, it holds that $\square{(f+g)}\geq\square f+\square g$;
  3. (3)
    if $s=(a_{1}\mathbbm{1}_{A_{1}},\dots,a_{n}\mathbbm{1}_{A_{n}})$, for $A_{i}\in\mathcal{A}_{i}$, then its sup-convolution $\square s$ is $\mathscr{F}$-measurable.
Note that our notion of a sup-convolution is dependent on the measure spaces $(E_{i},\mathscr{E}_{i},\mu_{i})$ and $(F,\mathscr{F},\nu)$, but it does not necessarily depend on an algebraic structure of said spaces.
For brevity we will hereafter refer to a “generalized sup-convolution” as just a sup-convolution. Recall that if $E_{1}=E_{2}=\mathbb{R}^{n}$ and $\mathcal{A}_{1}=\mathcal{A}_{2}=\mathcal{B}(\mathbb{R}^{n})$ the Borel $\sigma$-algebra, and $f=(f_{1},f_{2})\in\mathcal{F}(\mathcal{A}_{1}\times\mathcal{A}_{2})$
$\square f(z)\coloneqq\sup_{z=(1-t)x+ty}f_{1}^{1-t}(x)f_{2}^{t}(y)$
is a sup-convolution.
More generally, if $E_{1},E_{2},\dots,E_{n},F$ are Polish spaces, $t_{i}>0$, $\sum_{i=1}^{n}t_{i}=1$ and $\Phi:E\to F$ continuous, then
$\square f(z)\coloneqq\sup_{z=\Phi(x)}\prod_{i=1}^{n}f_{i}^{t_{i}}(x_{i})$
is a sup-convolution as well. This in a very particular case of the next result.
In what follows a set is said to be analytic if it is the continuous image of a Polish space.
Theorem 2.3.
Under the above conventions, given $\mathcal{A}=(\mathcal{A}_{1},\dots,\mathcal{A}_{n})$ with each $\mathcal{A}_{i}$ a sub-collection of analytic sets, for $\alpha\leq 1$ and $t=(t_{1},\dots,t_{n})$, with $t_{i}>0$ such that $\sum_{i=1}^{n}t_{i}=1$, let
$\mathcal{M}_{\alpha}^{(t)}(u_{i})=\left(\sum_{i=1}^{n}t_{i}u_{i}^{\alpha}\right)^{\frac{1}{\alpha}}$
and $C(z)=\{x:\Phi(x)=z\}$ for a continuous function $\Phi:E=E_{1},\times\cdots\times E_{n}\to F$. The map defined by
$\square f(z)\coloneqq\sup_{x\in C(z)}\mathcal{M}_{\alpha}^{(t)}(f_{i}(x_{i}))$
is a sup-convolution.
In the following proof (and later on) we call a function a “step” function if it takes only finitely many values.
Proof.
Take $f\leq g$ and $x\in C(z)$. Then $\mathcal{M}_{\alpha}^{(t)}(f_{i}(x_{i}))\leq\mathcal{M}_{\alpha}^{(t)}(g_{i}(x_{i}))\leq\square g(z)$. Taking supremum over $x\in C(z)$ shows that $\square$ is monotone.
We proceed to verifying superaddivity. Assume that $f=(f_{1},\dots,f_{n})$ and $g=(g_{1},\dots,g_{n})$, with $f_{i},g_{i}$ being step functions, such that if $g(x_{i})>0$, then $f_{i}(x_{i})\geq f_{i}(y_{i})$ for all $y_{i}\in E_{i}$. Now if $\square g(z)=0$, then the result is trivial from monotonicity. Otherwise, $g(z)>0$, and there exists [^5] Since $g_{i}$ are step functions $x\mapsto\mathcal{M}_{\alpha}^{(t)}(g_{i}(x_{i}))$ takes only finite many values. an $x\in C(z)$ such that $\square g(z)=\mathcal{M}_{\alpha}^{(t)}(g_{i}(x_{i}))>0$. For such an $x$, necessarily $g_{i}(x_{i})>0$ so that $f_{i}(x_{i})\geq f_{i}(y_{i})$ for all $y_{i}\in E_{i}$. It follows that
$\mathcal{M}_{\alpha}^{(t)}(f_{i}(x_{i}))=\square f(z).$
Consequently,
$\square f(z)+\square g(z)$ $=\mathcal{M}_{\alpha}^{(t)}(f_{i}(x_{i}))+\mathcal{M}_{\alpha}^{(t)}(g_{i}(x_{i}))$
$\leq\mathcal{M}_{\alpha}^{(t)}((f_{i}+g_{i})(x_{i}))$
$\leq\sup_{y\in C(z)}\mathcal{M}_{\alpha}^{(t)}((f_{i}+g_{i})(x_{i}))$
$=\square(f+g)(z),$
as required.
To establish measurability, given $s=(a_{1}\mathbbm{1}_{A_{1}},\dots,a_{n}\mathbbm{1}_{A_{n}})$, we have
$\square s(z)=\mathcal{M}_{\alpha}^{(t)}(a_{i})\mathbbm{1}_{\Phi(A)}(z).$
For each $i$, since $A_{i}$ is analytic, there exist Polish spaces $P_{i}$ and continuous functions $\psi_{i}$ such that $\psi_{i}(P_{i})=A_{i}$. Consider the Polish space $P=P_{1}\times\cdots\times P_{n}$, and define $\psi:P\to F$ by
$\psi(p)=\Phi(\psi_{1}(p_{1}),\dots,\psi_{n}(p_{n})).$
Then $\psi(P)=\Phi(A)$, so that $\Phi(A)$ is analytic, and universally measurable (see [26, Chapter 43]). It follows that $\square s$ is universally measurable as well. ∎
Our main tool in this paper is the following equivalence of geometric and functional inequalities.
Theorem 2.4.
Under the above conventions, given $\mathcal{A}=(\mathcal{A}_{1},\dots,\mathcal{A}_{n})$ with each $\mathcal{A}_{i}$ a sub-collection of analytic sets, for $\alpha\leq 1$ and $t=(t_{1},\dots,t_{n})$, with $t_{i}>0$ such that $\sum_{i=1}^{n}t_{i}=1$, a triple $(\nu,\mu,\square)$ and $p=(p_{1},\dots,p_{n})\in[0,\infty)^{n}$ satisfies
$\nu(\square a)\geq\mathcal{M}_{\alpha}^{(t)}(\|a_{i}\|_{p_{i}})$
for all functions $a$ of the form $a\coloneqq(a_{1}\mathbbm{1}_{A_{1}},\dots,a_{n}\mathbbm{1}_{A_{n}})$ for $a_{i}\in(0,\infty)$ and $A_{i}\in\mathcal{A}_{i}$ if and only if it satisfies for any collection of non-negative $f_{i}\in L_{p_{i}}(\mu_{i})\cap\mathcal{F}(\mathcal{A})$,
$\nu_{*}(\square f)\geq\mathcal{M}_{\alpha}^{(t)}(\|f_{i}\|_{p_{i}}),$
where $\nu_{*}$ denotes the lower integral of $\square f$.
We prepare for the proof with the next lemma.
Lemma 2.5.
For a set $U$ and function $f:U\to\mathbb{R}$ and $t\in\mathbb{R}$, set
$f^{-}\coloneqq\min\{f,t\}\ \ \ \hbox{ and }\ \ \ f^{+}\coloneqq\max\{f-t,0\}.$
If $f^{+}(x)>0$, then $f^{-}(x)\geq f^{-}(y)$ for all $y\in U$.
Proof.
For $x$ such that $f^{+}(x)>0$, $f(x)>t$, so that $f^{-}(x)=t\geq\min\{f(y),t\}=f^{-}(y)$. ∎
In what follows, for $p=(p_{1},\dots,p_{n})$ with $p_{i}>0$ we denote for brevity
$\|f\|_{p}\coloneqq\left(\|f_{1}\|_{p_{1}},\dots,\|f_{n}\|_{p_{n}}\right).$
Proof of Theorem 2.4.
Let us first observe that it is enough to prove the inequality for $f_{i}$ step functions. Define
$f_{i}^{(m)}\coloneqq\sum_{k=1}^{2^{2m}}2^{-m}\mathbbm{1}_{\left\{f_{i}\geq\frac{k}{2^{m}}\right\}},$
and set $f^{(m)}=(f_{1}^{(m)},\dots,f_{n}^{(m)})$. Note that $f_{i}\in\mathcal{F}(\mathcal{A}_{i})$ implies that $f_{i}^{(m)}\in\mathcal{F}(\mathcal{A}_{i})$ as well. Thus if the inequality holds for step functions, we have
$\nu_{*}\left(\square f\right)\geq\nu\left(\square f^{(m)}\right)\geq\mathcal{M}_{\alpha}^{(t)}\left(\|f^{(m)}\|_{p}\right),$
where the first inequality follows from the monotonicity of the sup-convolution, Definition 2.2 (1), and the fact that $f\geq f^{(m)}$. By the monotone convergence theorem, $\|f^{(m)}\|_{p}$ increases to $\|f\|_{p}$ and by the continuity of $\mathcal{M}_{\alpha}^{(t)}(\cdot)$, we get
$\mathcal{M}_{\alpha}^{(t)}(\|f\|_{p})=\lim_{m\to\infty}\mathcal{M}_{\alpha}^{(t)}(\|f^{(m)}\|_{p})\leq\nu(\square f).$
Thus we may assume that the $f_{i}$ take only finitely many values and prove the result through induction on the total number of non-zero values taken by the $f_{i}$. That is, if $N_{i}$ denotes the number of distinct non-zero values taken by $f_{i}$, we denote $N=\sum_{i=1}^{n}N_{i}$. The base case follows by hypothesis. Let us assume $N>n$, that the $\mathcal{F}(\mathcal{A})$-functional inequality holds for step functions who take strictly less than $N$ total values, and that all $f_{i}$ are non-zero step functions. For each $i=1,\dots,n$, we write
$f_{i}=\sum_{j=1}^{N_{i}}a_{ji}\mathbbm{1}_{A_{ji}},$
with $a_{ji}>0$, $A_{ji}\in\mathcal{A}_{i}$ and $A_{ji}\supseteq A_{(j+1)i}$, with $\sum_{i=1}^{n}N_{i}=N$. Define
$f_{1}^{-}=a_{11}\mathbbm{1}_{A_{11}}\hbox{ and }\ f_{1}^{+}=\sum_{j=2}^{N_{1}}a_{j1}\mathbbm{1}_{A_{j1}}.$
Note that $f_{1}^{-}$ and $f_{1}^{+}$ individually take strictly fewer values than $f_{1}$. Define
$\lambda=\frac{\|f^{-}_{1}\|_{p_{1}}}{\|f_{1}\|_{p_{1}}},$
so that $1-\lambda=\frac{\|f^{+}_{1}\|_{p_{1}}}{\|f_{1}\|_{p_{1}}}$. Note that $\lambda\in(0,1)$. Let
$f_{i}^{-}=\min\{f_{i},t_{i}\}\ \hbox{ and }\ f_{i}^{+}=\max\{f_{i}-t_{i},0\}.$
where $t_{i}$ is implicitly and uniquely defined through the equation
$\frac{\|f_{i}^{-}\|_{p_{i}}}{\|f_{i}\|_{p_{i}}}=\lambda$
because $t\mapsto\mu_{i}(\min\{f_{i},t\})$ is continuous and strictly increasing on $(0,\|f_{i}\|_{\infty})$.
Set $f^{+}=(f_{1}^{+},\dots,f_{n}^{+})$ and $f^{-}=(f_{1}^{-},\dots,f_{n}^{-})$. By Lemma 2.5 and Definition 2.2 Item (2)
$\square f\geq\square f^{-}+\square f^{+}.$
Integrating this inequality and applying the inductive hypothesis,
$\nu(\square f)$ $\geq\nu(\square f^{-})+\nu(\square f^{+})$
$\geq\mathcal{M}_{\alpha}^{(t)}(\|f^{-}_{i}\|_{p_{i}})+\mathcal{M}_{\alpha}^{(t)}(\|f^{+}_{i}\|_{p_{i}})$
$=\mathcal{M}_{\alpha}^{(t)}(\lambda\|f_{i}\|_{p_{i}})+\mathcal{M}_{\alpha}^{(t)}((1-\lambda)\|f_{i}\|_{p_{i}})$
$=\mathcal{M}_{\alpha}^{(t)}(\|f_{i}\|_{p_{i}}),$
as desired. ∎
Concerning operations in line with our concentrations are the articles [28, 29]. In [28] Gardner, Hug, and Weil conducted a systemic study of operations between sets in ${\mathbb{R}}^{n}$ with a particular emphasis on compact convex sets in ${\mathbb{R}}^{n}$. Among these operations are the Minkowski sum, and $L_{p}$-Minkowski sum (see Section 5), which we consider here. Additionally, the article [29] of Gardner and Kiderlin concerns what they called “point-wise, homogeneous, monotonic, and associative” between functions defined on ${\mathbb{R}}^{n}$, and they characterize such operations as representable by $40$ explicitly given types. In particular, [29, Section 11] concerns a characterization of the supremal convolution appearing in (6) on the class of logarithmically concave functions on ${\mathbb{R}}^{n}$, denoted $\text{LC}({\mathbb{R}}^{n})$. They show, [29, Theorem 11.1], that any operation $\star:\text{LC}({\mathbb{R}}^{n})\times\text{LC}({\mathbb{R}}^{n})\to\text{LC}({\mathbb{R}}^{n})$ verifying certain properties, in addition to the above ones, must arise as a sup-convolution. Though most of our focus in this work is in Euclidean space, the framework for our “generalized sup-convolution” is much more flexible, approaching inequalities in measure spaces, both geometric and functional. An analogous classification of the seems beyond reach currently.
3. Dimensional Functional Brunn-Minkowksi inequalities in measure spaces
In this section we will extend several Brunn-Minkowski type inequalities from recent literature to functional versions. To start, as mentioned in the introduction the following result was conjectured by Gardner and Zvavitch in [30] and confirmed by Eskenazis and Moschidis in [24].
Theorem A (Eskenazis & Moschidis, [24]).
Let $\gamma$ be the standard Gaussian probability measure on ${\mathbb{R}}^{n}$ with density given by $\frac{d\gamma}{dx}=(2\pi)^{-\frac{n}{2}}e^{-\frac{|x|^{2}}{2}}$, For $t\in(0,1)$ and $A,B$ origin symmetric convex bodies in $\mathbb{R}^{n}$, it holds that
$\gamma((1-t)A+tB)\geq\mathcal{M}^{(t)}_{\frac{!}{n}}(\gamma(A),\gamma(B)).$
We prove the following Borell-Brascamp-Lieb type inequality holds for the Gaussian measure (namely Theorem 1.1) when restricting ourselves to even unimodal functions; this constitutes an improvement of the Prékopa-Leindler inequality (3).
Proof of Theorem 1.1.
We prove the result first when $\alpha=-\frac{1}{n}$. Take $\mathcal{A}=(\mathcal{A}_{1},\mathcal{A}_{2})$ for $\mathcal{A}_{i}$ the cone of all symmetric convex bodies, and $\Phi:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}^{n}$ to be $\Phi(x,y)=(1-t)x+ty$ and $C(z)=\{(x,y):\Phi(x,y)=z\}$. According to Theorem 2.3,
$f\square g(z)\coloneqq\sup_{(1-t)x+ty=z}\mathcal{M}^{(t)}_{-\frac{1}{n}}(f(x),g(y))$
is a sup-convolution. Thus, in view of Theorem 2.4, our result holds if it holds for scaled indicator functions. Suppose that $f=a\mathbbm{1}_{A}$ and $g=b\mathbbm{1}_{B}$ for $A$ and $B$ convex bodies. For the remainder of the proof, we will always assume that $f$ and $g$ are of the above form. Observe that
$\gamma(f\square g)$ $=\mathcal{M}_{-\frac{1}{n}}^{(t)}(a,b)\gamma((1-t)A+tB)$
$\geq\mathcal{M}_{-\frac{1}{n}}^{(t)}(a,b)\mathcal{M}_{\frac{1}{n}}^{(t)}(\gamma(A),\gamma(B))$
$\geq\min\{a\gamma(A),b\gamma(B)\},$
completing the proof for $\alpha=-\frac{1}{n}$.
Assume next that $\alpha>-\frac{1}{n}$. Then, by applying Holder’s inequality (9), we obtain
$h((1-t)x+ty)$ $\geq\mathcal{M}_{\alpha}^{(t)}(f(x),g(y))$
$=\mathcal{M}_{\alpha}^{(t)}\left(\frac{f(x)}{\gamma(f)}\gamma(f),\frac{g(y)}{\gamma(g)}\gamma(g)\right)$
$\geq\mathcal{M}_{-\frac{1}{n}}^{(t)}\left(\frac{f(x)}{\gamma(f)},\frac{g(y)}{\gamma(g)}\right)\mathcal{M}_{r}^{(t)}(\gamma(f),\gamma(g)).$
Applying the $\alpha=-\frac{1}{n}$ result to this inequality completes the proof of the theorem. ∎
This can be extended by the same method, using the following result.
Theorem B (Cordero-Erausquin & Rotem, [20]).
Let $w$ be an increasing function such that $s\mapsto w(e^{s})$ is convex, and $\mu$ the measure on $\mathbb{R}^{n}$ with density $\frac{d\mu}{dx}=e^{-w(|x|)}$. Then for symmetric convex bodies $A$ and $B$
$\mu((1-t)A+tB)\geq\mathcal{M}_{\frac{1}{n}}^{(t)}(\mu(A),\mu(B)).$
Theorem 3.1.
For $t\in(0,1)$, $\alpha\geq-\frac{1}{n}$ and $\mu$ a measure on $\mathbb{R}^{n}$ given by a density $\frac{d\mu}{dx}=e^{-w(|x|)}$ for $w$ increasing with $s\mapsto w(e^{s})$ convex, then if
$h((1-t)x+ty)\geq\mathcal{M}_{\alpha}^{(t)}(f(x),g(y))$
holds for $h,f,g:\mathbb{R}^{n}\to[0,\infty)$, with $f,g$ symmetric unimodal functions, we have
$\mu(h)\geq\mathcal{M}_{r}^{(t)}(\mu(f),\mu(g))$
for $r=\frac{\alpha}{1+n\alpha}$.
Proof.
The proof follows by the same arguments as that of Theorem 1.1 while applying Theorem B. ∎
Theorem C (Aishwarya & Rotem, [2]).
Let $V:\mathbb{R}^{n}\to[0,\infty)$ a $p$-homogeneous convex function for $1<p<\infty$, and let $\nu$ be a measure with density $\frac{d\nu}{dx}$ proportional to $e^{-V}$. Let $K_{0}$ and $K_{1}\subseteq\mathbb{R}^{n}$ be star bodies. Then for $0<t<1$,
$\nu((1-t)K_{0}+tK_{1})\geq\mathcal{M}^{(t)}_{\frac{p-1}{pn}}(\nu(K_{0}),\nu(K_{1}))$
Here $K$ is a star body if $K$ is Borel and satisfies $\lambda K\subseteq K$ or $\lambda\in[0,1]$ and $V$ a $p$-homogeneous function means that $V(\lambda x)=\lambda^{p}V(x)$ for $\lambda\geq 0$ and $x\in\mathbb{R}^{n}$. We consider a function $f:\mathbb{R}^{n}\to\mathbb{R}$ to be star-unimodal, if $t\in\mathbb{R}$ implies $\{f\geq t\}$ is a star body.
Theorem 3.2.
Let $V\colon{\mathbb{R}}^{n}\to[0,\infty)$ be $s$-homogeneous convex function for some $1<p<\infty$, and let $\nu$ be the measure with density $\frac{d\nu}{dx}=e^{-V}$, and let $t\in(0,1)$ and $\alpha\geq\frac{1-s}{sn}$. If $h,f,g:\mathbb{R}^{n}\to[0,\infty)$ with $f,g$ star-unimodal, are a triple of functions satisfying
$h((1-t)x+ty)\geq\mathcal{M}_{\alpha}^{(t)}(f(x),g(y)),$
then
$\nu(h)\geq\mathcal{M}_{r}^{(t)}(\nu(f),\nu(g)),$
where $r=\frac{\alpha(s-1)}{\alpha sn+s-1}$.
Proof.
Again, the methods of Theorem 1.1 can be applied to $\mathcal{A}_{i}$ the collection of star bodies and applying Theorem C. ∎
Finally we give a functional version of the recent work of Cordero-Erausquin and Eskenazis, [19, Corollary 2].
Theorem D (Cordero-Erausquin & Eskenazis, [19]).
Let $w\colon[0,\infty)\to(-\infty,\infty]$ be an increasing function such that $t\mapsto w(e^{t})$ is a convex function, and suppose that $\mu$ is the measure on ${\mathbb{R}}^{n}$ given by $d\mu(x)=e^{-w(|x|)}dx$. Suppose additionally that $C\subset{\mathbb{R}}^{n}$ is an origin symmetric convex set and $\varphi\colon C\to[0,\infty)$ is even and concave. Given $\beta>0$, the measure defined by $d\nu_{\beta}(x)=\varphi(x)^{\beta}d\mu(x)$, which is supported in $C$, verifies the following inequality: for any $t\in(0,1)$ and and pair of origin symmetric convex bodies $K,L\subset{\mathbb{R}}^{n}$, it holds that
$\nu_{\beta}((1-t)K+tL)\geq\mathcal{M}_{\frac{1}{\beta+n}}^{(t)}(\nu_{\beta}(K),\nu_{\beta}(L)).$
Thanks to Theorem 2.3 and Theorem 2.4, we now show that Theorem D above is equivalent to a functional version involving even unimodal functions (note that the direction from functions to sets is obvious).
Theorem 3.3.
Let $\beta>0$, $\alpha\geq-\frac{1}{\beta+n}$, and functions $h,f,g:\mathbb{R}^{n}\to[0,\infty)$, such that $f,g$ are unimodal and even, and satisfy
$h((1-t)x+ty)\geq\mathcal{M}_{\alpha}^{(t)}(f(x),g(y))$
for $x,y\in\mathbb{R}^{n}$. Then, for $\nu_{\beta}$ satisfying the assumptions of Theorem D,
$\nu_{\beta}(h)\geq\mathcal{M}_{r}^{(t)}\left(\nu_{\beta}(f),\nu_{\beta}(g)\right),$
for $r=\frac{\alpha(\beta+n)}{\alpha+n+\beta}$.
Proof.
Proving the result $\alpha=-\frac{1}{\beta+n}$ and $\mathcal{A}=(\mathcal{A}_{1},\mathcal{A}_{2})$ for $\mathcal{A}_{i}$ the cone of all symmetric convex bodies, and $\Phi:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}^{n}$ to be $\Phi(x,y)=(1-t)x+ty$ and $C(z)=\{(x,y):\Phi(x,y)=z\}$, and applying Theorem 2.3 and then extending to $\alpha>-\frac{1}{\beta+n}$ by Hölder’s inequality (9)as in the Proof of Theorem 1.1, the result follows.
4. Applications to Barthe’s Reverse Brascamp-Lieb Inequality and Schneider’s Conjecture
Here we focus on applications of our results to inequalities concerning the Lebesgue measure on ${\mathbb{R}}^{n}$. In the first subsection, we will treat Barthe’s reverse Brascamp-Lieb inequality proven in [6, 7], while in the second subsection we will show how our results can be used to understand an older conjecture of Schneider from [65] concerning generalized Brunn-Minkowski-type inequalities.
4.1. Barthe’s Reverse Brascamp-Lieb inequality
The celebrated Brascamp–Lieb inequality from [15], asserts that for $m,n\in\mathbb{N}$, $n_{i}\in\mathbb{N}$ and $c_{i}>0$ such that $c_{1}n_{1}+\ldots+c_{m}n_{m}=n$, $B_{i}:\mathbb{R}^{n}\to\mathbb{R}^{n_{i}}$ surjective linear maps, and $f_{i}\in L^{1}(\mathbb{R}^{n_{i}})$ non-negative, the following inequality is valid
$\int_{\mathbb{R}^{n}}\prod_{i=1}^{m}f_{i}(B_{i}x)^{c_{i}}dx\leq D\prod_{i=1}^{m}\left(\int_{\mathbb{R}^{n_{i}}}f_{i}\right)^{c_{i}},$
where the constant can be computed by considering just centered Gaussians (see [40] and also [8] for the characterization of $D<+\infty$). In [6, 7], Barthe proved the following reverse form of Brascamp–Lieb inequality.
Theorem E (Barthe).
Let $m,n$ be non-negative integers, $(c_{i})_{1}^{m}$ be a connections of positive real numbers and $(n_{i})_{1}^{m}$ be a collection of non-negative integers smaller than $n$ such that
$\sum_{i=1}^{m}c_{i}n_{i}=n.$
For $i=1,\dots,m$ let $B_{i}$ be a linear surjective map from $\mathbb{R}^{n}$ onto $\mathbb{R}^{n_{i}}$ and set $f=(f_{1},\dots,f_{m})$, with $f_{i}\colon{\mathbb{R}}^{n_{i}}\to[0,\infty)$. If
$\bigcap_{i=1}^{m}\text{ker}B_{i}=\{0\}.$
then the functional
$I(f)=\int_{\mathbb{R}^{n}}^{*}\sup\left\{\prod_{i=1}^{m}f_{i}^{c_{i}}(x_{i})\colon\sum_{i=1}^{m}c_{i}B_{i}^{*}x_{i}=z\hbox{ and }x_{i}\in\mathbb{R}^{n_{i}}\right\}dz.$
satisfies the inequality
(10) $I(f)\geq E\prod_{i=1}^{m}\left(\int_{\mathbb{R}^{n_{i}}}f_{i}\right)^{c_{i}},$
where
(11) $\begin{split}E&=\inf\left\{\frac{I(g)}{\prod_{i=1}^{m}\left(\int_{\mathbb{R}^{n_{i}}}g_{i}\right)^{c_{i}}}\colon g_{i}\hbox{ centered Gaussian on }\mathbb{R}^{n_{i}},i=1,\dots,m\right\}\\ &=\sqrt{\inf\left\{\frac{\det\left(\sum_{i=1}^{m}c_{i}B_{i}^{*}M_{i}B_{i}\right)}{\prod_{i=1}^{m}(\det(M_{i}))^{c_{i}}}\colon M_{i}\in S^{+}(\mathbb{R}^{n_{i}}),i=1,\dots,m\right\}},\end{split}$
and $S^{+}({\mathbb{R}}^{n_{i}})$ denotes the collection of symmetric positive definite $n_{i}\times n_{i}$ matrices.
Set $c\coloneqq\sum_{i=1}^{n}c_{i}$, and for each $i=1,\dots,m,$, let $t_{i}=\frac{c_{i}}{c}$ and $\tilde{f}_{i}=f_{i}^{c}$. Then we can rewrite the inequality (10) in the form
$\int_{\mathbb{R}^{n}}^{*}\sup\left\{\prod_{i=1}^{m}\tilde{f}_{i}^{t_{i}}(x_{i})\colon\sum_{i=1}^{m}c_{i}B_{i}^{*}x_{i}=z\hbox{ and }x_{i}\in\mathbb{R}^{n_{i}}\right\}dz\geq E\prod_{i=1}^{m}\left(\int_{\mathbb{R}^{n_{i}}}\tilde{f}_{i}^{\frac{1}{c}}\right)^{ct_{i}}.$
Next, by writing $f$ in place of $\tilde{f}$ above, we have the following equivalent formulation of Barthe’s reverse Brascamp-Lieb inequality (10):
$\int_{\mathbb{R}^{n}}^{*}\sup\left\{\prod_{i=1}^{m}{f}_{i}^{t_{i}}(x_{i});\sum_{i=1}^{m}c_{i}B_{i}^{*}x_{i}=z\hbox{ and }x_{i}\in\mathbb{R}^{n_{i}}\right\}dz\geq E\prod_{i=1}^{m}\left(\|{f}_{i}\|_{\frac{1}{c}}\right)^{t_{i}}.$
Theorem 1.4 will follow by combining the next results.
Proposition 4.1.
With the definitions of this section, the map
$\square f(z)\coloneqq\sup_{x\in C(z)}\prod_{i=1}^{n}f_{i}^{t_{i}}(x_{i}),$
where
$C(z)\coloneqq\left\{x\in\mathbb{R}^{n_{1}}\times\cdots\times{\mathbb{R}}^{n_{m}}:\sum_{i=1}^{m}c_{i}B_{i}^{*}x_{i}=z\right\},$
defines a sup-convolution.
Proof.
The goal is the verify the condition of the Definition 2.2. Notice that monotonicity is obvious, while the superadditivity condition follows from the fact that the geometric mean satisfies a reverse Minkowski inequality. Finally, measurability follows from the fact that $\Phi(x)=\sum_{i=1}^{m}c_{i}B_{i}^{*}x_{i}$ is a continuous function. ∎
Corollary 4.2.
The reverse Brascamp–Lieb inequality (10) is equivalent to the following geometric inequality:
(12) $\left|\sum_{i=1}^{m}c_{i}B_{i}^{*}A_{i}\right|\geq E\ \mathcal{M}_{0}^{(c)}\left(|A_{i}|_{n_{i}}\right).$
for sets $A_{i}\subset{\mathbb{R}}^{n_{i}}$.
Proof.
We start with the forward direction. By taking $f_{i}=\mathbbm{1}_{A_{i}}$, we have $\sup_{x\in C(z)}\mathcal{M}_{0}^{(c)}(f_{i}(x_{i}))=\mathbbm{1}_{\sum_{i=1}^{m}c_{i}B_{i}^{*}A_{i}}$, whereby inequality (10) implies
$\int_{{\mathbb{R}}^{n}}\sup_{x\in C(z)}\mathcal{M}_{0}^{(c)}(f_{i}(x_{i}))dz=\left|\sum_{i=1}^{m}c_{i}B_{i}^{*}A_{i}\right|\geq E\prod_{i=1}^{m}\int_{{\mathbb{R}}^{n_{i}{}}}f_{i}=E\ \mathcal{M}_{0}^{(c)}(|A_{i}|_{n_{i}}).$
Conversely, consider the sup-convolution of Proposition 4.1 with $t_{i}\coloneqq\frac{c_{i}}{\|c\|_{1}}$ for $\|c\|_{1}\coloneqq\sum_{i=1}^{m}c_{i}$ and note that by Theorem 2.4 and the homogeneity of the inequality, in order to prove
(13) $\int_{{\mathbb{R}}^{n}}\square f\geq E\ \mathcal{M}_{0}^{(t)}\left(\|f_{i}\|_{{\|c\|_{1}}^{-1}}\right),$
it suffices to check the inequality for $f_{i}=\mathbbm{1}_{A_{i}}$, in which case the inequality is just (12). Now apply (13) to the function $f_{i}=f_{i}^{\|c\|_{1}}$, to obtain
$\int_{{\mathbb{R}}^{n}}\sup_{x\in C(z)}\prod_{i=1}^{m}f^{c_{i}}_{i}(x_{i})dz\geq E\ \prod_{i=1}^{m}\left(\int_{{\mathbb{R}}^{n_{i}}}f_{i}^{\frac{\|c\|_{1}}{\|c\|_{1}}}\right)^{t_{i}\|c\|_{1}}=E\ \prod_{i=1}^{m}\left(\int_{{\mathbb{R}}^{n_{i}}}f_{i}\right)^{c_{i}},$
and we recover (10), as desired. ∎
Remark 4.3.
Such an approach characterizes the finiteness of the constant (11) associated to the reverse Brascamp-Lieb, when one wants to restrict to some class of functions $\mathcal{F}(\mathcal{A})$, to be equivalent with the finiteness of the geometric constant in the inequality (12) for the restricted class $\mathcal{A}$, as it can be seen by repeating verbatim the above proof for this class.
4.2. A Functional Version of Schneider’s Conjecture
In [65] Schneider considered the following generalization of Minkowski difference (called the $m$-order difference body): given a convex body $K\subset\mathbb{R}^{n}$, and $n\geq 2,m\geq 1$, define the $m$-th order difference body of $K$ via
$D^{m}(K)=\{x=(x_{1},\dots,x_{m})\in(\mathbb{R}^{n})^{m}\colon K\cap(x_{1}+K)\cdots\cap(x_{m}+K)\neq\emptyset\}.$
In [65], Schneider established the following variant of the celebrated Rogers-Shephard inequality:
(14) $|D^{m}(K)|_{nm}|K|^{-m}\leq\binom{n(m+1)}{n}$,
with equality only for $K$ a simplex in ${\mathbb{R}}^{n}$. Schneider additionally proved the following generalization of the Brunn-Minkowski inequality for $n=2$ and $m$ arbitrary. For any convex body $K\subset{\mathbb{R}}^{2}$, it holds that
(15) $|D^{m}(K)|_{2m}|K|_{2}^{-m}\geq(m+1)^{2},$
with equality if and only if $K$ is centrally symmetric. Note that for $n\geq 1$ and $m=1$, the above reduces to the Brunn-Minkowski inequality applied to KKK-K.
Various extensions of known geometric operators were inspired by the construction appearing in [65]; this theory been rapidly developing (see [60, 31, 35, 37, 34, 36, 32]). In particular a functional version of the inequality (14) was proven in [34], while a measure theoretic variant was established in [60].
The following fascinating question was left open in [65].
Conjecture 4.4 (Schneider’s conjecture (1970)).
For $n\geq 3$ and $m\geq 2$, is $|D^{m}(K)|_{nm}|K|^{-m}$ minimized by ellipsoids? In order words, is it necessarily the case that
$\min_{K\subset{\mathbb{R}}^{n}}|D^{m}(K)|_{nm}|K|^{-m}=|D^{m}(B)|_{nm}|B|^{-m}$
whenever $n\geq 3,m\geq 2$? Here $B\subset{\mathbb{R}}^{n}$ is the $\ell_{2}$ unit ball.
It is worth noting that this question has an intimate connection with a long standing conjecture of Petty appearing in [57], but we do no elaborate on this connection here. Very recently, in [32], Conjecture 4.4 was proven with $D^{m}K$ replaced by $(D^{m}K)^{\circ}$, the polar set, where the maximizers were shown to be ellipsoids. Thus, pairing this result with the bounds for the volume product, Conjecture 4.4 is known to hold up to an absolute constant. Additionally, the authors proved a functional version of this dual inequality using the functional Blaschke-Santalo inequality and Barthe’s reverse Brascamp-Lieb inequality.
Proposition 4.5.
Let $m,n$ be positive integers and $\mathcal{A}$ be the collection of convex bodies in ${\mathbb{R}}^{n}$. Consider $\bar{f}=(f,\dots,f)$, with $f\in\mathcal{F}({\mathcal{A}})$ being non-negative. Then the map $\square_{m}\bar{f}\colon({\mathbb{R}}^{n})^{m}\to[0,\infty)$ defined by
$\square_{m}\bar{f}(z)\coloneqq\left(\sup_{x\in\mathbb{R}^{n}}\prod_{i=0}^{m}f(x-z_{i})\right)^{\frac{1}{m+1}}$
is a sup-convolution.
Proof.
The monotonicity is clear from the definition of the supremum, which verifies (1) from Definition 2.2. Suppose that $\bar{f}=(f,\dots,f),\bar{g}=(g,\dots,g)$ are step functions for which $f(x_{i})\leq f(y_{i})$ for all $y_{i}\in{\mathbb{R}}^{n}$ whenever $x_{i}\in{\mathbb{R}}^{n}$ is such that $g(x_{i})>0$. We need to prove that $\square_{m}(f+g)\leq\square_{m}f+\square_{m}g$. Should $g\equiv 0$, then the result is obvious. Suppose that $g\not\equiv 0$, and that $z$ is such that $\square g(z)>0$, and choose $\varepsilon>0$ sufficiently small such that $\square g(z)-\varepsilon>0$. Next, select $x\in{\mathbb{R}}^{n}$ such that
$g(x)g(x-z_{1})\cdots g(x-z_{m})\geq\square_{m}\bar{g}(z)-\varepsilon.$
Necessarily, for each $i$, this we obtain $f(x-z_{i})>0$, whence $f(x-z_{i})\geq f(y-z_{i})$ holds for all $y$; hence,
$f(x)f(x-z_{1})\cdots f(x-z_{m})=\square_{m}\bar{f}(z).$
Letting $z_{0}=0$, we obtain
$\square_{m}\bar{f}(z)+\square_{m}\bar{g}(z)$ $\leq g(z)g(x-z_{1})\cdots g(x-z_{m})+f(x)f(x-z_{1})\cdots f(x-z_{m})$
$\leq\prod_{i=0}^{m}\left(g(x-z_{i})+f(x-z_{i})\right)+\varepsilon$
$\leq\sup_{y\in{\mathbb{R}}^{n}}\prod_{i=0}^{m}\left(g(y-z_{i})+f(y-z_{i})\right)+\varepsilon$
$=\square_{m}(f+g)(z)+\varepsilon.$
Definition 2.2 (2) follows from the inequality $\prod_{i=0}^{m}a_{i}+\prod_{i=0}^{m}b_{i}\leq\prod_{i=0}^{m}(a_{i}+b_{i})$ for $a_{i},b_{i}\geq 0$. Taking $\varepsilon\to 0$ completes the proof. Finally, the condition (3) follows from the measurability of $D^{m}(K)$.
Corollary 4.6.
Let $m\geq 2,n\geq 3$. The geometric inequality in Conjecture 4.4 is equivalent to the functional inequality: given $\bar{f}=(f,\dots,f)$, for $f\colon{\mathbb{R}}^{n}\to[0,\infty)$ unimodal, it holds that
(16) $\int_{(\mathbb{R}^{n})^{m}}\square_{m}\bar{f}(z)\geq c(m,n)\left(\int_{\mathbb{R}^{n}}f^{\frac{1}{m}}\right)^{m},$
where $c(m,n)=|D^{m}(B)|_{nm}|B|^{m}$ and $B$ is the $\ell_{2}$ unit ball.
Proof.
To prove that the functional inequality implies Schneider’s geometric conjecture, take $f=\mathbbm{1}_{K}$, so that $\square f(z)=\mathbbm{1}_{D^{m}(K)}$, the integral inequality is then
$|D^{m}(K)|_{nm}\geq c(m,n)|K|^{m},$
which is exactly Schneider’s conjectured inequality. If Schneider’s conjectured inequality holds, then the inequality holds for $f$ the indicator of a convex set, noting that the inequality is homogeneous we have the result for $f$ a scaled step function. Interpreting $\left(\int_{\mathbb{R}^{n}}f^{\frac{1}{m}}\right)^{m}=\mathcal{M}_{1}^{(1/2)}(\|f\|_{\frac{1}{m}})$ where $\mathcal{M}_{1}^{(1/2)}(u)=u$ as it is the mean of one value, and applying Theorem 2.4 the result follows. ∎
We mention in passing that the following operation considered in [34] can be confirmed by analogous methods to be a sup-convolution as well.
Definition 4.7.
For $\mathcal{A}$ the space of convex bodies and $f_{i}:\mathbb{R}^{n}\to[0,\infty)$ belonging to $\mathcal{F}(\mathcal{A})$, define $\square{f}:(\mathbb{R}^{n})^{m}\to[0,\infty)$
$\square f(z)\coloneqq\sup_{x\in\mathbb{R}^{n}}\prod_{i=0}^{m}f_{i}(x-z_{i}),$
where $z_{0}\coloneqq 0$ for notational brevity.
We note that Schneider’s conjecture is confirmed for convex sets $K\subseteq\mathbb{R}^{n}$ when $n=2$ and $m\geq 1$ is arbitrary, hence (16) is a valid functional inequality for $n\leq 2$ by (15), which constitutes a generalization of the Prékopa-Leindler inequality (3) when $n=2$ for the function
$\square_{1}f(z)=\sup_{z=x-y}f(x)f(y).$
Theorem 4.8.
Let $n=2$, $m\geq 1$. Then, for any unimodal function $f\colon{\mathbb{R}}^{n}\to[0,\infty)$, we have
$\int_{(\mathbb{R}^{2})^{m}}\square_{m}f(z)\geq(m+1)^{2}\left(\int_{\mathbb{R}^{2}}f^{\frac{1}{m}}\right)^{m}.$
Moreover, this functional inequality is equivalent to (15).
Proof.
To prove that the functional inequality implies the geometric inequality (15), it is enough choose $f=\mathbbm{1}_{K}$, whereby $\square_{m}f=\mathbbm{1}_{D^{m}(K)}$. Conversely, combining the inequality (15) with Theorem 2.4 completes the proof. ∎
5. Functional $L_{p}$-Brunn-Minkowski inequalities
In this section we investigate how our methods can be used to provide functional inequalities in the $L_{p}$-Brunn-Minkowski theory. In particular, we develop a functional equivalent of the celebrated $\log$-Brunn-Minkowski conjecture from [14].
This theory has been pivotal in many advancements in convex geometry and affine differential geometry over the last thirty years. The literature is extensive, so we direct the reader to the articles [25, 42, 43, 52, 44, 45, 46, 47, 49, 48, 50], which demonstrate the versatility of this subject and its connection to other fields. All of the above concern the case of $p\geq 1$.
On the other hand, very recently the study of the case $p\in[0,1]$ has gained much momentum; see the survey [12] of Böröczky for a detailed history, and related results and conjectures, and their connections to other fields. In particular, we would like to direct the reader’s attention to the works [14, 54, 41, 63, 64, 18, 33, 16].
We remind the reader that we take a convex body to be a compact convex set with non-empty interior. For convex bodies containing the origin in their interior, $K_{0}$ and $K_{1}$, $t\in(0,1)$ and $p\in[0,\infty)$ one defines the $L_{p}$-Minkowski convex combination of $K_{0}$ with $K_{1}$, denoted $(1-t)\cdot K_{0}+_{p}t\cdot K_{1}$, via the Wülff shape of $p$-th average of the support functions of the $K_{i}$. More explicitly, for a set $A\subseteq\mathbb{R}^{n}$ define $h_{A}:\mathbb{R}^{n}\to\mathbb{R}$ by
$h_{A}(z)\coloneqq\sup_{x\in A}\langle x,z\rangle.$
Let $(1-t)\cdot K_{0}+_{p}t\cdot K_{1}$ be defined as
(17) $(1-t)\cdot K_{0}+_{p}t\cdot K_{1}:=\bigcap_{\theta\in\mathbb{S}^{n-1}}\left\{x\in\mathbb{R}^{n}:\langle x,\theta\rangle\leq\mathcal{M}_{p}^{(t)}\left(h_{K_{0}}(\theta),h_{K_{1}}(\theta)\right)\right\}.$
Note that when $p=1$ the $L_{p}$-Minkowski sum amounts to the classical Minkowski addition $(1-t)K_{0}+tK_{1}$.
For $p=0$ we shorten the notation and write
$K_{0}^{1-t}K_{1}^{t}\coloneqq(1-t)\cdot K_{0}+_{0}t\cdot K_{1}.$
In the case $p>1$, and $q=\frac{p}{p-1}$, this definition extends to Borel sets in [51] in the following way: given Borel measurable sets $A,B\subset{\mathbb{R}}^{n}$ and $t\in(0,1)$, define
$(1-t)A+_{p}tB$ $=\bigcup_{\lambda\in[0,1]}\left\{(1-t)^{\frac{1}{p}}(1-\lambda)^{\frac{1}{q}}A+t^{\frac{1}{p}}\lambda^{\frac{1}{q}}B\right\}$
$=\left\{z:\exists(a,b,\lambda)\in A\times B\times[0,1]\text{ with }\ z=(1-t)^{\frac{1}{p}}(1-\lambda)^{\frac{1}{q}}a+t^{\frac{1}{p}}\lambda^{\frac{1}{q}}b\right\}.$
As proved in [51, Lemma 1-2] this interpretation of the $L_{p}$-Minkowski sum agrees with (17) when $p>1$ and when $A,B$ are convex bodies containing the origin in their interiors. Therein, the authors established the following general form of the $L_{p}$-Brunn-Minkowski inequality for Borel sets for $p>1$:
(18) $|(1-t)A+_{p}tB|\geq\mathcal{M}_{\frac{p}{n}}^{(t)}(|A|,|B|).$
Since $(1-t)A+tB\subseteq(1-t)A+_{p}tB$ for $p>1$ (a non-trivial inclusion proved in [51, Lemma 3]), inequality (18) follows from the classical Brunn-Minkowski inequality (1) and a homogeneity argument.
For $p\in[0,1)$, the following conjecture – that has deeply influenced the research in convex geometry in recent years – was put forth by Böröczky, Lutwak, Yang and Zhang in [14]:
Conjecture 5.1 ($L_{p}$-Brunn-Minkowski Conjecture [14]).
Let $0\leq p<1$. Given any pair of origin symmetric convex bodies $K,L\subset{\mathbb{R}}^{n}$ and $t\in(0,1)$, does it necessarily follow that
$|(1-t)K+_{p}tL|\geq\mathcal{M}_{\frac{p}{n}}^{(t)}(|K|,|L|)?$
It should be noted that the inequality appearing in Conjecture 5.1 dramatically improves the classical Brunn-Minkowski inequality. The case $p=0$ is known as the $\log$-Brunn-Minkowski conjecture. Work towards this conjecture and related problems are rapidly developing and we do not attempt to cover all of the related papers and results. Instead we choose to focus on a few important cases for which this conjecture was verified: for $n=2$ in [14], locally near a Euclidean ball in [18], and for unconditional convex bodies in [63] when $p=0$ and for $0<p<1$ in [53]. Further it was verified in [64] that the $\log$-Brunn-Minkowski inequality for volume implies the same inequality with volume replaced by any even $\log$-concave measure on ${\mathbb{R}}^{n}$; a similar result for $0<p<1$ was established in [41]. In [33], a general treatment of local $L_{p}$-Brunn-Minkowski inequalities were treated, and in [16], Conjecture 5.1 was verified for to hold for all $p\in(p_{0},1)$ for some $p_{0}$ depending on the dimension.
In the following subsections we will demonstrate how our methods can be used to prove functional versions of geometric inequalities appearing in the $L_{p}$-Brunn-Minkowski theory. We also provide a functional equivalent of the geometric inequality appearing in Conjecture 5.1.
5.1. The case $p\in[0,1)$
Here we move to the case $p\in[0,1)$. We begin with a definition.
Definition 5.2.
Let $0\leq p\leq 1$. Let $\mathcal{A}$ be the cone of all origin symmetric convex bodies in $\mathbb{R}^{n}$, so that $\mathcal{F}(\mathcal{A})$ is the collection of all even unimodal functions. For $f=(f_{0},f_{1})$, with $f_{i}\in\mathcal{F}(\mathcal{A})$, define the sup-convolution
(19) $\square_{p}f(z)\coloneqq f_{0}\square_{p}f_{1}(z)\coloneqq\sup_{(x_{0},x_{1})\in C(z;f)}\mathcal{M}_{-\frac{p}{n}}^{(t)}(f_{0}(x_{0}),f_{1}(x_{1})),$
where
$C_{p}(z;f)\coloneqq\left\{(x_{0},x_{1}):z\in(1-t)\cdot\{f_{0}\geq f_{0}(x_{0})\}+_{p}t\cdot\{f_{1}\geq f_{1}(x_{1})\}\right\}.$
The next theorem assets that $\square_{p}$ is indeed a sup-convolution, in the sense of Definition 2.2, when $p\in[0,1]$.
Theorem 5.3.
Let $p\in[0,1]$. For $\mathcal{A}$ the cone of origin symmetric convex bodies, let $f=(f_{0},f_{1})$, with $f_{i}\in\mathcal{F}(\mathcal{A})$. Then the map $\square_{p}f$ defined in (19) is a sup-convolution.
Proof.
We first prove monotonicity. Let $f=(f_{0},f_{1}),g=(g_{0},g_{1})$ with $f_{i},g_{i}\in\mathcal{F}(\mathcal{A})$ satisfying $f_{i}\leq g_{i}$. Note that if $f_{i}$ is the zero function, $\square_{p}f(z)$ is identically zero and there is nothing to prove. Now suppose that $f_{i},g_{i}\in\mathcal{F}(\mathcal{A})$, are such that $f_{i}\leq g_{i}$ for all $i$. Fix $x=(x_{0},x_{1})\in C_{p}(z;f)$ such that $f_{0}(x_{0})f_{1}(x_{1})>0$. By the condition $f_{i}\leq g_{i}$, we have
$\{f_{i}\geq f_{i}(x_{i})\}\subseteq\{g_{i}\geq f_{i}(x_{i})\}.$
Since $\{g_{i}\geq f_{i}(x_{i})\}\in\mathcal{A}$, it is necessarily compact, and hence there exists $y_{i}\in\{g_{i}\geq f_{i}(x_{i})\}$ such that $g_{i}(y_{i})\leq g_{i}(x)$ for all $x$ such that $g_{i}(x)\geq f_{i}(x_{i})$. Thus
$\{g_{i}\geq g_{i}(y_{i})\}=\{g_{i}\geq f_{i}(x_{i})\}\supseteq\{f_{i}\geq f_{i}(x_{i})\}.$
Thus, $h_{\{g_{i}\geq g_{i}(y_{i})\}}\geq h_{\{f_{i}\geq f_{i}(x_{i})\}}$, and so we have that
$z\in(1-t)\cdot\{f_{0}\geq f_{0}(x_{0})\}+_{p}t\cdot\{f_{1}\geq f_{1}(x_{1})\}\subseteq(1-t)\cdot\{g_{0}\geq g_{0}(y_{0})\}+_{p}t\cdot\{g_{1}\geq g_{1}(y_{1})\}.$
Consequently, $(y_{0},y_{1})\in C_{p}(z;g)$, and hence
$\mathcal{M}_{-\frac{p}{n}}^{(t)}(f_{0}(x_{0}),f_{1}(x_{1}))\leq\mathcal{M}_{-\frac{p}{n}}^{(t)}(g_{0}(y_{0}),g_{1}(y_{1}))\leq\sup_{w\in C(z;g)}\mathcal{M}_{-\frac{p}{n}}^{(t)}(g_{0}(w_{0}),g_{1}(w_{1}))=\square_{p}g(z).$
Taking the sup over all $x\in C_{p}(z,f)$ completes the proof of monotonicity.
Now suppose that $f_{i}$ and $g_{i}$ are step functions and are such that $g_{i}(x_{i})>0$ implies that $f_{i}(x_{i})\geq f_{i}(y_{i})$ for all $y_{i}\in\mathbb{R}^{n}$. We wish to show that $\square_{p}(f+g)\geq\square_{p}f+\square_{p}g$. If $\square_{p}g(z)=0$, then $\square_{p}(f+g)(z)\geq\square_{p}f(z)$ by monotonicity. Thus we can assume that $\square_{p}g(z)>0$. Given $\varepsilon>0$ there exists $x=(x_{0},x_{1})\in C(z;g)$ such that
$\mathcal{M}_{-\frac{p}{n}}^{(t)}(g_{0}(x_{0}),g_{1}(x_{1}))>\square_{p}g(z)-\varepsilon$
and $g_{i}(x_{i})>0$ for all $i$. Since $f_{i}(x_{i})\geq f_{i}(y_{i})$ for all $y_{i}$ (by our assumption), we get
$\mathcal{M}_{-\frac{p}{n}}^{(t)}(f_{0}(x_{0}),f_{1}(x_{1}))\geq\mathcal{M}_{-\frac{p}{n}}^{(t)}(f_{0}(y_{0}),f_{1}(y_{1})).$
Taking the suprema of the right hand side of the above equation over $y\in C_{p}(z;f)$ gives
$\mathcal{M}_{-\frac{p}{n}}^{(t)}(f_{0}(x_{0}),f_{1}(x_{1}))\geq\square_{p}f(z).$
However, since $x\in C_{p}(z;g)$ we have by the hypothesis on $f$ and $g$,
$\{g_{i}\geq g_{i}(x_{i})\}\subseteq\{g_{i}>0\}\subseteq\{f_{i}\geq f_{i}(x_{i})\},$
so that
$z\in(1-t)\cdot\{g_{0}\geq g_{0}(x_{0})\}+_{p}t\cdot\{g_{1}\geq g_{1}(x_{1})\}\subseteq(1-t)\cdot\{f_{0}\geq f_{0}(x_{0})\}+_{p}t\cdot\{f_{1}\geq f_{1}(x_{1})\},$
whence $x\in C_{p}(z;f)$. Thus,
$\mathcal{M}_{-\frac{p}{n}}^{(t)}(f_{0}(x_{0}),f_{1}(x_{1}))=\square_{p}f(z).$
Applying the reverse triangle inequality, $\mathcal{M}_{-\frac{p}{n}}^{(t)}(u_{0},u_{1})+\mathcal{M}_{-\frac{p}{n}}^{(t)}(v_{0},v_{1})\leq\mathcal{M}_{-\frac{p}{n}}^{(t)}(u_{0}+v_{0},u_{1}+v_{1})$, we have
$\square_{p}f(z)+\square_{p}g(z)$ $\leq\mathcal{M}_{-\frac{p}{n}}^{(t)}(f_{0}(x_{0}),f_{1}(x_{1}))+\mathcal{M}_{-\frac{p}{n}}^{(t)}(g_{0}(x_{0}),g_{1}(x_{1}))+\varepsilon$
$\leq\mathcal{M}_{-\frac{p}{n}}^{(t)}((f_{0}+g_{0})(x_{0}),(f_{1}+g_{1})(x_{1}))+\varepsilon.$
To complete the proof, it remains to see that $x\in C_{p}(z;f+g)$ so that the suprema can be taken, and then $\varepsilon\to 0$. But this is immediate, as $\{g_{i}\geq g_{i}(x_{i})\}\subseteq\{f_{i}\geq f_{i}(x_{i})\}=\{f_{i}=f_{i}(x_{i})\}$, we infer that $\{g_{i}\geq g_{i}(x_{i})\}\subseteq\{f_{i}+g_{i}\geq(f_{i}+g_{i})(x_{i})\}$; and we have
$z\in(1-t)\cdot\{g_{0}\geq g_{0}(x_{0})\}$ $+_{p}t\cdot\{g_{1}\geq g_{1}(x_{1})\}$
$\subseteq(1-t)\cdot\{(f_{0}+g_{0})\geq(f_{0}+g_{0})(x_{0})\}+_{p}t\cdot\{(f_{1}+g_{1})\geq(f_{1}+g_{1})(x_{1})\}.$
For $f_{0}=\mathbbm{1}_{A}$ and $f_{1}=\mathbbm{1}_{B}$ we claim that
$\square_{p}f(z)=\mathbbm{1}_{(1-t)\cdot A+_{p}t\cdot B}(z).$
$\square_{p}f(z)=1$ if there exists $a\in A$ and $b\in B$ such that $(a,b)\in C_{p}(z;f)$ and is 0 otherwise, but $(a,b)\in C_{p}(z,f)$ is equivalent to
$z\in(1-t)\cdot\{f_{0}\geq f_{0}(a)\}+_{p}t\cdot\{f_{1}\geq f_{1}(b)\}=(1-t)\cdot\{f_{0}\geq 1\}+_{p}t\cdot\{f_{1}\geq 1\}=(1-t)\cdot A+_{p}t\cdot B.$
Thus, the measurability of $\square_{p}f$ in this case follows from the fact that $(1-t)\cdot A+_{p}t\cdot B$ is a convex set. ∎
The next statement establishes an equivalent functional version of the $L_{p}$-Brunn-Minkowski inequality.
Theorem 5.4.
For $p\in[0,1)$, the $L_{p}$-Brunn-Minkowski inequality
$|(1-t)\cdot A+_{p}t\cdot B|\geq\mathcal{M}_{\frac{p}{n}}^{(t)}(|A|,|B|)$
holding for all origin symmetric convex bodies $A$ and $B$ in ${\mathbb{R}}^{n}$ is equivalent to the following functional inequality: given $\alpha\geq-\frac{p}{n}$, $t\in(0,1)$, and a triple of functions $h,f,g\colon{\mathbb{R}}^{n}\to[0,\infty)$, with $f,g$ even and unimodal, satisfying
$h(z)\geq\mathcal{M}_{\alpha}^{(t)}(f(x),g(y)),$
where
$(x,y)\in C_{p}(z;f,g)\coloneqq\{(u,v):z\in(1-t)\cdot\{f\geq f(u)\}+_{p}t\cdot\{g\geq g(v)\},$
then
$\int_{{\mathbb{R}}^{n}}h\geq\mathcal{M}_{r}^{(t)}\left(\int_{{\mathbb{R}}^{n}}f,\int_{{\mathbb{R}}^{n}}g\right)$
for $r=\frac{\alpha p}{p+\alpha n}$.
Proof.
If the set theoretic result holds, by Theorem 5.3 the operation $f\square_{p}g$ is a sup-convolution, so by Theorem 2.4 is suffices to prove the functional inequality when $f=a\mathbbm{1}_{A}$ and $g=b\mathbbm{1}_{B}$, where $A$ and $B$ are origin symmetric convex bodies, and $a,b>0$. In this case
$h(z)\geq f\square_{p}g(z)=\mathcal{M}_{-\frac{p}{n}}^{(t)}(a,b)\mathbbm{1}_{(1-t)\cdot A+_{p}t\cdot B}.$
Integrating both side and applying Hölder’s inequality (9) and the $L_{p}$-Brunn-Minkowski inequality, we have
$\int_{{\mathbb{R}}^{n}}h\geq\int_{{\mathbb{R}}^{n}}f\square_{p}g$ $=\mathcal{M}_{-\frac{p}{n}}^{(t)}(a,b)|(1-t)\cdot A+_{p}t\cdot B|$
$\geq\mathcal{M}_{-\frac{p}{n}}^{(t)}(a,b)\mathcal{M}_{\frac{p}{n}}^{(t)}(|A|,|B|)$
$=\mathcal{M}_{-\infty}^{(t)}(a|A|,b|B|).$
This yields the functional inequality when $\alpha=-\frac{p}{n}$. When $\alpha>-\frac{p}{n}$, we have by assumption, and then by Hölder’s inequality (9), that
$h(z)$ $\geq\sup_{(x,y)\in C(z;f,g)}\mathcal{M}_{\alpha}^{(t)}\left(\frac{f(x)\int_{{\mathbb{R}}^{n}}f}{\int_{{\mathbb{R}}^{n}}f},\frac{g(y)\int_{{\mathbb{R}}^{n}}g}{\int_{{\mathbb{R}}^{n}}g}\right)$
$\geq\sup_{(x,y)\in C(z;f,g)}\mathcal{M}_{-\frac{p}{n}}^{(t)}\left(\frac{f(x)}{\int_{{\mathbb{R}}^{n}}f},\frac{g(y)}{\int_{{\mathbb{R}}^{n}}g}\right)\mathcal{M}_{r}^{(t)}\left(\int_{{\mathbb{R}}^{n}}f,\int_{{\mathbb{R}}^{n}}g\right).$
Now observing the scaling invariance of $C(z;f,g)$, that is that $C(z;\lambda f,tg)=C(z;f,g)$ for $\lambda,t>0$ we see that the above inequality can be written as
$h(z)\geq\frac{f}{\int_{{\mathbb{R}}^{n}}f}\square_{p}\frac{g}{\int_{{\mathbb{R}}^{n}}g}(z)\mathcal{M}_{r}^{(t)}\left(\int_{{\mathbb{R}}^{n}}f,\int_{{\mathbb{R}}^{n}}g\right).$
Integrating, we conclude that
$\int_{{\mathbb{R}}^{n}}h$ $\geq\mathcal{M}_{r}^{(t)}\left(\int_{{\mathbb{R}}^{n}}f,\int_{{\mathbb{R}}^{n}}g\right)\int_{{\mathbb{R}}^{n}}\left(\frac{f}{\int_{{\mathbb{R}}^{n}}f}\square_{p}\frac{g}{\int_{{\mathbb{R}}^{n}}g}\right)$
$\geq\mathcal{M}_{r}^{(t)}\left(\int_{{\mathbb{R}}^{n}}f,\int_{{\mathbb{R}}^{n}}g\right)\mathcal{M}_{-\infty}\left(\int_{{\mathbb{R}}^{n}}\frac{f}{\int_{{\mathbb{R}}^{n}}f},\int_{{\mathbb{R}}^{n}}\frac{g}{\int_{{\mathbb{R}}^{n}}g}\right)$
$=\mathcal{M}_{r}^{(t)}\left(\int_{{\mathbb{R}}^{n}}f,\int_{{\mathbb{R}}^{n}}g\right).$
For the converse, if there exists $\alpha\geq-\frac{p}{n}$ such that the functional inequality holds, with $A$ and $B$ origin symmetric convex bodies and let $f=a\mathbbm{1}_{A}$, $g=b\mathbbm{1}_{B}$. Applying the functional inequality gives
$\mathcal{M}_{\alpha}^{(t)}(a,b)|(1-t)\cdot A+_{p}t\cdot B|\geq\mathcal{M}_{r}^{(t)}(a|A|,b|B|)=\mathcal{M}_{\alpha}^{(t)}(a,b)\mathcal{M}_{\frac{p}{n}}^{(t)}(|A|,|B|),$
where the last equality gives the result by taking $a=|A|^{\frac{p}{\alpha n}}$ $b=|B|^{\frac{p}{\alpha n}}$, chosen to satisfy equality in Hölder. ∎
Again, in cases where geometric inequalities are already verified, with Theorem 5.4 in hand, one obtains new $L_{p}$-Borell-Brascamp-Lieb type inequalities for $p<1$. For example we use the following result from [16].
Theorem F (Chen, Huang, Li, & Liu, [16]).
There exists a $p_{0}\in(0,1)$, depending on the dimension, such that for all $p\in(p_{0},1)$ and for any pair of origin-symmetric convex bodies $K,L\subset{\mathbb{R}}^{n}$ and any $t\in[0,1]$, it holds that
$|(1-t)\cdot K+_{p}t\cdot L|\geq\mathcal{M}_{\frac{p}{n}}^{(t)}(|K|,|L|).$
Theorem 5.5.
Let $p_{0}$ be as in Theorem F and let $p\in(p_{0},1)$. Let $\alpha\geq-\frac{p}{n}$, and $h,f,g\colon{\mathbb{R}}^{n}\to[0,\infty)$ be a triple of functions, with $f,g$ being even and unimodal, such that
$h(z)\geq\mathcal{M}_{\alpha}^{(t)}(f(x),g(y))$
for
$(x,y)\in C_{p}(z;f,g)\coloneqq\{(u,v):z\in(1-t)\cdot\{f\geq f(u)\}+_{p}t\cdot\{g\geq g(v)\}.$
Then the following $L_{p}$-Borell-Brascamp-Lieb inequality holds:
$\int_{{\mathbb{R}}^{n}}h\geq\mathcal{M}^{(t)}_{r}\left(\int_{{\mathbb{R}}^{n}}f,\int_{{\mathbb{R}}^{n}}g\right),$
where $r=\frac{p\alpha}{p+\alpha n}$.
Next, we continue with the case of $p=0$. For starters, we take note in the next results that the $\log$-Brunn-Minkowski conjecture enjoys the following self-improvement results, which respectively appear in [64] and [41].
Theorem G (Saroglou, [64]).
If $|A^{1-t}B^{t}|\geq|A|^{1-t}|B|^{t}$ holds for all origin symmetric convex bodies $A,B\subset{\mathbb{R}}^{n}$ and all $t\in[0,1]$, then
$\mu(A^{1-t}B^{t})\geq\mu^{1-t}(A)\mu^{t}(B)$
whenever $A,B$ are origin symmetric convex bodies and $\mu$ is an even log-concave measure.
Theorem H (Livshyts, Marsiglietti, Nayar and Zvavitch, [41]).
For $\mu$ an even log-concave measure such that
$\mu(A^{1-t}B^{t})\geq\mu^{1-t}(A)\mu^{t}(B)$
holds for all $t\in[0,1]$ and $A,B\subset{\mathbb{R}}^{n}$ origin symmetric convex bodies, then for $p>0$
$\mu((1-t)\cdot A+_{p}t\cdot B)\geq\mathcal{M}_{\frac{p}{n}}^{(t)}(\mu(A),\mu(B))$
holds for all $t\in[0,1]$ and $A,B$ symmetric convex bodies.
As consequence of these results and Theorem 5.4, we obtain the following equivalent form of Conjecture 5.1.
Corollary 5.6.
The log-Brunn-Minkowski conjecture is true if and only if its functional extension holds. Moreover the log-Brunn-Minkowski inequality implies that the $L_{p}$-Borell-Brascamp-Lieb inequalities hold for all unimodal symmetric functions with respect to even log-concave measures.
In other words, if the $\log$-Brunn-Minkowski inequality holds, then for every $\alpha\geq-\frac{p}{n}$ and $h,f,g:\mathbb{R}^{n}\to[0,\infty)$ with $f$ and $g$ even and unimodal, such that
$h(z)\geq\mathcal{M}_{\alpha}^{(t)}(f(x),g(y))$
for
$(x,y)\in C_{p}(z;f,g)\coloneqq\{(u,v):z\in(1-t)\cdot\{f\geq f(u)\}+_{p}t\cdot\{g\geq g(v)\},$
then for a symmetric log-concave measure $\mu$,
$\mu(h)\geq\mathcal{M}_{r}^{(t)}\left(\mu(f),\mu(g)\right),$
where $r=\frac{p\alpha}{p+n\alpha}$.
The next result can be deduced from the articles [63, 64].
Theorem I (Saroglou, [63, 64]).
Let $\mu$ be a $\log$-concave measure with an unconditional density on Then, for any pair of unconditional convex bodies $K,L\subset{\mathbb{R}}^{n}$ and any $t\in[0,1]$,
$\mu(K^{1-t}L^{t})\geq\mathcal{M}_{0}^{(t)}(\mu(K),\mu(B)).$
Combining Theorem 5.4, with Theorem I and Corollary 5.6, we get Theorem 1.3
5.2. The case $p>1$
In this subsection we describe the $p>1$ a Borell-Brascamp-Lieb inequality that can be derived for $L_{p}$-Minkowski sums. The ideas are similar but easier. These results should be compared to [61, 62]. Here $q$ is the Hölder conjugate to $p$ satisfying $q=\frac{p}{p-1}.$
Definition 5.7.
Let $p>1$, $\alpha\in[-\infty,\infty]$ and Borel measurable functions $f_{0},f_{1}\colon{\mathbb{R}}^{n}\to[0,\infty)$ and fixed $t\in[0,1]$. For $f=(f_{0},f_{1})$, define
$\square_{p,\alpha}f(z)=\sup_{(x,y)\in C_{p}(z)}\mathcal{M}_{\alpha}^{(t)}(f_{0}(x),f_{1}(y)),$
where
$C_{p}(z):=\left\{(x,y)\colon\text{ there exists }\lambda\in[0,1]\text{ with }z=(1-t)^{\frac{1}{p}}(1-\lambda)^{\frac{1}{q}}x+t^{\frac{1}{p}}\lambda^{\frac{1}{q}}y\right\}.$
Note that in the $p>1$ case, for Borel $A$ and $B$ then
(20) $(1-t)\cdot A+_{p}t\cdot B=\Phi([0,1]\times A\times B)$
under the map $\Phi(\lambda,x,y)=(1-t)^{\frac{1}{p}}(1-\lambda)^{\frac{1}{q}}x+t^{\frac{1}{p}}\lambda^{\frac{1}{q}}y$. We have the following result.
Proposition 5.8.
Let $p>1$, $\alpha\geq-\frac{p}{n}$, $t\in(0,1)$, and $f=(f_{0},f_{1})$, with $f_{0},f_{1}\colon{\mathbb{R}}^{n}\to[0,\infty)$ Borel measurable functions, then if
$h(z)\geq\mathcal{M}_{\alpha}^{(t)}(f_{0}(x),f_{1}(y))$
for $(x,y)\in C_{p}(z),$
$\int_{\mathbb{R}^{n}}h\geq\mathcal{M}_{\alpha^{\prime}}^{(t)}\left(\int_{\mathbb{R}^{n}}f_{0},\int_{\mathbb{R}^{n}}f_{1}\right),$
where $\alpha^{\prime}=\frac{\alpha p}{p+\alpha n}$.
Proof.
The proof follows analogously to the results in Section 3. Start with the case $\alpha=-\frac{p}{n}$, where $\square_{p,\alpha}f$ is checked to be a sup-convolution through the argument of Theorem 2.3. Indeed, monotonicity is obvious, measurability follows from (20) which realizes $(1-t)\cdot A+_{p}t\cdot B$ as a the continuous image of a Polish space (namely $0,1]\times A\times B$ when $A$ and $B$ are assumed to be Borel), and the reverse triangle inequality follows from the same argument as before; for fixed $z$, find $(x,y)$ that $\mathcal{M}_{-\frac{p}{n}}^{(t)}(g_{0}(x),g_{1}(y))$ approximates $\square g(z)$ and by the hypothesis $\mathcal{M}_{-\frac{p}{n}}^{(t)}(f_{0}(x),f_{1}(y))=\square f(z)$, and let the reverse triangle inequality for $\mathcal{M}_{-\frac{p}{n}}^{(t)}$ conclude. Using Theorem 2.4 we obtain the result for $\alpha=-\frac{p}{n}$. For $\alpha>-\frac{p}{n}$, homogeneity again allows us to conclude. ∎
6. An Abstract Approach to the Equivalence of Geometric and Functional Inequalities
In this section the goal is to create an abstract framework for the equivalence between a generic inequality of Brunn-Minkowski type and its functional counterpart, with respect to more general means. We show that a relevant sup-convolution for the Borell-Ehrhard inequality does unfortunately not yield an equivalence of the functional and geometric version of the inequality. We close with a Borell-Brascamp-Lieb type inequality for nilpotent Lie groups.
In what follows we will consider measure spaces $(E_{1},\mu_{1}),\ldots,(E_{n},\mu_{n})$ and $(F,\nu)$, and set as before $E=E_{1}\times\cdots\times E_{n}$ and $\mu=\mu_{1}\otimes\cdots\otimes\mu_{n}$.
Definition 6.1.
Let $m_{i}\geq\mu_{i}(E_{i})$, for $i=1,\ldots,n$, $I:=\prod_{i=1}^{n}[0,m_{i}]$ and $J:=\prod_{i=1}^{n}[0,m_{i}^{2}]$. We consider three generalized means
$\mathcal{W},\mathcal{M}:I\to[0,\infty),\text{ and }\mathcal{N}:J\to[0,\infty),$
that vanish whenever a coordinate is 0 and satisfy the following properties:
  1. (1)
    for $u,v\in I$ such that $u+v\in I$,
    $\mathcal{M}(u+v)\geq\mathcal{M}(u)+\mathcal{M}(v);$
  2. (2)
    for $u\in J$ and $\lambda\in(0,1)$,
    $\mathcal{N}(\lambda u)+\mathcal{N}((1-\lambda)u)\geq\mathcal{N}(u);$
  3. (3)
    for $u,v\in I$, a “Hölder” inequality is satisfied in the sense
    $\mathcal{W}(u)\mathcal{M}(v)\geq\mathcal{N}(uv),$
    where $(uv)_{i}=u_{i}v_{i}$.
Remark 6.2.
Usually one just considers $E_{i}=F$ for each $i$, while $m_{i}\in\{1,+\infty\}$. An important example, when $m_{i}=+\infty$ is supplied by
$\mathcal{W}(u)\coloneqq\left(\sum_{i=1}^{n}t_{i}u_{i}^{\alpha}\right)^{\frac{1}{\alpha}}$
for $t_{i}>0$ and $\alpha\geq-1$,
$\mathcal{M}(u)\coloneqq\left(\sum_{i=1}^{n}t_{i}u_{i}^{p}\right)^{\frac{1}{p}}$
for $p\in[-\alpha,1]$, and
$\mathcal{N}(u)=\left(\sum_{i=1}^{n}t_{i}u_{i}^{r}\right)^{\frac{1}{r}}.$
Definition 6.3.
To each point $z\in F$, we assume that there exists a subset of $C(z)\subset E$. For any $f=(f_{1},\dots,f_{n})$, with $f_{i}:E_{i}\to(0,m_{i}]$, we define the $C$-sup-convolution, $\square f:F\to\mathbb{R}$ by
$\square f(z)\coloneqq\sup_{x\in C(z)}\mathcal{M}(f(x)).$
Proposition 6.4.
Let $\mathcal{M}$ be as above and $C(z):=\Phi^{-1}(z)$ for a continuous function $\Phi:E\to F$. Then the $C$-sup-convolution is a sup-convolution in the sense of Definition 2.2.
Proof.
For $x\in C(z)$, since by assumption $f_{i}(x_{i})\leq g_{i}(x_{i})$ we have by monotonicity of means,
$\mathcal{M}(f_{i}(x_{i}))\leq\mathcal{M}(g_{i}(x_{i})).$
Taking supremums concludes the monotonicity of $\square$.
To prove the superadditivity and measurability properties, we proceed exactly as in Theorem 2.3. For the former we employ now property (1) from Definition 6.1. ∎
Definition 6.5.
For $A=(A_{1},\dots,A_{n})$ with $A_{i}\subseteq E_{i}$ define
$m(A)\coloneqq m(A_{1},\dots,A_{n})\coloneq\{z\in F:\exists x\in(A_{1}\times\cdots\times A_{n})\cap C(z)\}.$
Definition 6.6.
For $\mathcal{A}=(\mathcal{A}_{1},\dots,\mathcal{A}_{n})$ where each $\mathcal{A}_{i}$ is a class of subsets on $E_{i}$ , we say that $(E,\mu)$ and $(F,\nu)$ satisfy a $\mathcal{A}$-restricted geometric inequality if $A=(A_{1},\ldots,A_{n})\in\mathcal{A}$ implies that
$\nu(m(A))\geq\mathcal{W}(\mu_{i}(A_{i})).$
Definition 6.7.
The measure spaces $(E,\mu)$ and $(F,\nu)$ satisfy an $\mathcal{F}(\mathcal{A})$-functional inequality if $f_{i}\in\mathcal{F}(\mathcal{A}_{i})$ implies that
$\nu_{*}(\square f)\geq\mathcal{N}(\mu_{i}(f_{i})),$
where $\nu_{*}(f)\coloneqq sup_{\{s\leq f:simple\}}\nu(s)$.
The main theorem of this section is the next result.
Theorem 6.8.
If the measure spaces $(E,\mu)$ and $(F,\nu)$ satisfy an $\mathcal{A}$-restricted geometric inequality, then they satisfy a $\mathcal{F}(\mathcal{A})$-functional inequality as well. Conversely, assume that the means $\mathcal{W},\mathcal{M}$ and $\mathcal{N}$ satisfy the following “duality” relation: For all $u\in I$,
(21) $\sup_{(a_{1},\ldots,a_{n})\in I}\frac{\mathcal{N}(a_{1}u_{1},\ldots,a_{n}u_{n})}{\mathcal{M}(a_{1},\ldots,a_{n})}=\mathcal{W}(u).$
Then the $\mathcal{F}(A)$-functional inequality implies the $\mathcal{A}$-restricted geometric inequality.
Proof.
The proof follows as in Theorem 2.3. One reduces to step functions via monotone convergence. Thus we assume that the $f_{i}$ take only finitely many values and prove the result through induction on the total number of non-zero values taken by the $f_{i}$. We write as in the proof of Theorem 2.4
$f_{i}=\sum_{j_{i}=1}^{n_{i}}a_{j_{i}}\mathbbm{1}_{A_{j_{i}}},$
and a decomposition into $f_{i}^{+}$ and $f_{i}^{-}$, which allow,
$\nu(\square f)$ $\geq\nu(\square f^{-})+\nu(\square f^{+})$
$\geq\mathcal{N}(\mu_{i}(f_{i}^{-}))+\mathcal{N}(\mu_{i}(f_{i}^{+}))$
$=\mathcal{N}(\lambda\mu_{i}(f_{i}))+\mathcal{N}((1-\lambda)\mu_{i}(f_{i}))$
$\geq\mathcal{N}(\mu_{i}(f_{i})),$
where we have used the definition of $\lambda$ in the equality and Item (2) to conclude. To finish, for $f_{i}=a_{i}\mathbbm{1}_{A_{i}}$, $\square f=\mathcal{M}(a_{i})\mathbbm{1}_{m(A)}$ so that
$\mu(\square f)=\mathcal{M}(a_{i})\mu(m(A))\geq\mathcal{M}(a_{i})\mathcal{W}(\mu_{i}(A_{i}))\geq\mathcal{N}(a_{i}\mu_{i}(A_{i}))=\mathcal{N}(\mu_{i}(f_{i})).$
For the converse, consider $A_{i}\in\mathcal{A}_{i}$ and $a_{i}\in(0,m_{i}]$. The $\mathcal{F}(\mathcal{A})$-functional inequality for the functions $f_{i}:=a_{i}\mathbbm{1}_{A_{i}}$ implies
$\nu(m(A))\mathcal{M}(a_{1},\ldots,a_{n})\geq\mathcal{N}(a_{1}\mu_{1}(A_{1}),\ldots,a_{n}\mu_{n}(A_{n}))$
and thus
$\nu(m(A))\geq\sup_{a_{i}\in(0,m_{i}]}\frac{\mathcal{N}\left(a_{1}\mu_{1}(A_{1}),\ldots,a_{n}\mu_{n}(A_{n})\right)}{\mathcal{M}(a_{1},\ldots,a_{n})}=\mathcal{W}(\mu_{1}(A_{1}),\ldots,\mu_{n}(A_{n})).$
We take this moment to consider an example of a Brunn-Minkowski-type inequality that doesn’t fit into our framework.
Example 6.9.
For $t\in{\mathbb{R}}$, define
$\Phi(t):=\Phi_{\gamma}(t)=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{t}e^{-\frac{t^{2}}{2}}dt.$
For each $t\in(0,1)$, define the mean $\mathcal{M}_{\gamma}^{(t)}=\mathcal{M}_{\gamma}$ on $0,1]^{2}$ via
$\mathcal{M}_{\gamma}(a):=\Phi\left(t\Phi^{-1}(a_{1})+(1-t)\Phi^{-1}(a_{2})\right).$
The Borell-Ehrhard inequality ([21, 22, 23, 11]) asserts that
$t\Phi^{-1}(\gamma(A))+(1-t)\Phi^{-1}(\gamma(B))\leq\Phi^{-1}\left(\gamma(tA+(1-t)B)\right).$
For $A=[\Phi^{-1}(a_{1}),\Phi^{-1}(a_{2})]$ and $B=[\Phi^{-1}(b_{1}),\Phi^{-1}(b_{2})]$ it gives
$t\Phi^{-1}(a_{2}-a_{1})+(1-t)\Phi^{-1}(b_{2}-b_{1})\leq\Phi^{-1}\left(\Phi(t\Phi^{-1}(a_{2})+(1-t)\Phi^{-1}(b_{2}))-\Phi(t\Phi^{-1}(a_{1})+(1-t)\Phi^{-1}(b_{1}))\right),$
or equivalently,
$\Phi\left(t\Phi^{-1}(a_{2}-a_{1})+(1-t)\Phi^{-1}(b_{2}-b_{1})\right)\leq\Phi(t\Phi^{-1}(a_{2})+(1-t)\Phi^{-1}(b_{2}))-\Phi(t\Phi^{-1}(a_{1})+(1-t)\Phi^{-1}(b_{1})).$
In particular, the mean $\mathcal{M}_{\gamma}$ satisfies Property (1) of Definition 6.1 and the associated operation is indeed a sup-convolution, when $C(z)=\{(x,y):\ z=tx+(1-t)y\}$.
On the other hand, let $f=a_{1}\mathbbm{1}_{(-\infty,\Phi^{-1}(\lambda)]}$ and $g=a_{2}\mathbbm{1}_{(-\infty,\Phi^{-1}(\lambda)]}$, then for $h=\mathcal{M}_{\gamma}(a)\mathbbm{1}_{(-\infty,\Phi^{-1}(\lambda)]}$ the functional version of Borell-Ehrhard inequality gives
$\Phi^{-1}\left(\lambda\mathcal{M}_{\gamma}(a)\right)\geq t\Phi^{-1}(\lambda a_{1})+(1-t)\Phi^{-1}(\lambda a_{2}),$
or in other words,
$\lambda\mathcal{M}_{\gamma}(a)\geq\mathcal{M}_{\gamma}(\lambda a).$
This implies that the mean $\mathcal{M}_{\gamma}$ does not satisfy property (2) of Definition 6.1. This shows that the Borell-Ehrhard inequality does not fit the present framework.
Numerous Brunn-Minkowski inequalities have been suggested in the setting of groups. In particular, for the Heisenberg group in [56, 39, 5], and extentions thereof [9, 67, 58]. Our machinery shows Tao’s Prékopa-Leindler inequality [67, Theorem 4] for nilpotent groups is equivalent to its set theoretic formulation, and consequently to the following. Given a simply connected nilpotent Lie group $G$ and subsets $A,B$, we define their product set to be
$A\cdot B=\{x\cdot y\colon x\in A,y\in B\},$
where $\cdot$ denotes the group product operation. For more details nilpotent Lie groups see [67, 58].
Theorem 6.10.
Let $G$ be a simply connected nilpotent Lie group of (topological) dimension $d$ and with a Haar measure $\mu$. Then, for $\alpha\in[-\frac{1}{d},1]$ and $t\in(0,1)$, if $f,g,h\colon G\to[0,\infty)$ are $\mu$-measurable functions such that
$h(x\cdot y)\geq\mathcal{M}_{\alpha}^{(t)}(f(x),g(y))$
holds for all $x,y$, then
$\int_{G}h\ d\mu\geq\mathcal{M}_{\alpha^{\prime}}^{(t)}\left(\frac{\int_{G}f\ d\mu}{(1-t)^{d}},\frac{\int_{G}g\ d\mu}{t^{d}}\right),$
where $\alpha^{\prime}=\frac{\alpha}{1+\alpha d}$.
Proof.
By [67, Theorem 3] or [58, Theorem 1.1], we see that
(22) $\mu^{\frac{1}{d}}(A\cdot B)\geq\mu^{\frac{1}{d}}(A)+\mu^{\frac{1}{d}}(B).$
Taking
$f\square g(z)=\sup_{(x,y)\in C(z)}\mathcal{M}_{\alpha}^{(t)}(f(x),g(y))$
with $C(z)=\{(x,y):x\cdot y=z\}$, we have a sup-convolution, as $(x,y)\mapsto x\cdot y$ is continuous, Theorem 2.3 applies. By Theorem 2.4 it suffices to prove the result when $f=a\mathbbm{1}_{A}$ and $g=b\mathbbm{1}_{B}$ which reduces, after applying (22), to the following
$\int_{G}f\square_{\alpha}gd\mu$ $=\mathcal{M}_{\alpha}^{(t)}(a,b)\mu(A\cdot B)$
$\geq\mathcal{M}_{\alpha}^{(t)}(a,b)\left(\mu^{\frac{1}{d}}(A)+\mu^{\frac{1}{d}}(B)\right)^{d}$
$=\mathcal{M}_{\alpha}^{(t)}(a,b)\mathcal{M}_{\frac{1}{d}}^{(t)}\left(\frac{\mu(A)}{(1-t)^{d}},\frac{\mu(B)}{t^{d}}\right)$
$\geq\mathcal{M}_{\alpha^{\prime}}^{(t)}\left(\frac{\mu(f)}{(1-t)^{d}},\frac{\mu(g)}{t^{d}}\right)$
with $\frac{1}{\alpha}+d=\frac{1}{\alpha^{\prime}}$ so that $\alpha^{\prime}=\frac{\alpha}{1+\alpha d}$. ∎
Note for $\alpha>1$, the relevant operator $\square$ does not define a sup-convolution, but mirroring the Euclidean case, one can obtain for instance for any $\alpha\geq-\frac{1}{d},$
$h(x\cdot y)$ $\geq\mathcal{M}_{\alpha}^{(t)}(f(x),g(y))$
$=\mathcal{M}_{\alpha}^{(t)}\left(\frac{f(x)}{\mu(f)}\mu(f),\frac{g(y)}{\mu(g)}\mu(g)\right)$
$\geq\mathcal{M}_{-\frac{1}{d}}^{(t)}\left(\frac{f(x)}{\mu(f)},\frac{g(y)}{\mu(g)}\right)\mathcal{M}_{\alpha^{\prime}}(\mu(f),\mu(g)),$
where $\frac{1}{\alpha}=\frac{1}{\alpha^{\prime}}-d$, so that $\alpha^{\prime}=\frac{\alpha}{1+\alpha d}$. Applying the $\alpha=-\frac{1}{d}$ result gives
$\int_{G}h\ d\mu\geq\mathcal{M}_{-\infty}^{(t)}\left(\frac{1}{(1-t)^{d}},\frac{1}{t^{d}}\right)\mathcal{M}_{\alpha^{\prime}}^{(t)}(\mu(f),\mu(g))=\left(\frac{1}{\max\{1-t,t\}}\right)^{d}\mathcal{M}_{\alpha^{\prime}}^{(t)}\left(\int_{G}f\ d\mu,\int_{G}g\ d\mu\right),$
an analogous but weaker inequality in the case that $\alpha\leq 1$.
Acknowledgements
This project was inspired in part by Gautam Aishwarya’s enthusiasm for Theorem 1.1 and its proof. The second author thanks Aishwarya and HIM in Bonn Germany for hosting these discussions during the Dual Trimester Program “Boolean Analysis in Computer Science”. He thanks Tomasz Tkocz for encouragement and helpful discussions as well.
Malliaris and Roberto are supported by the FP2M federation (CNRS FR 2036); this work received support from the Graduate School EUR-MINT (State support managed by the National Research Agency for Future Investments program bearing the reference ANR-18-EURE-0023).
Melbourne is supported by CONAHCYT Grant CBF2023-2024-3907.
Roysdon is supported by US NSF Grant DMS-2452384.